### 31 Dec 2023 v1

# Stability for the 2-D plane Poiseuille flow in finite channel

Abstract.
In this paper, we study the stability for 2-D plane Poiseuille flow $(1-y^{2},0)$ in a channel $\mathbb{T}\times(-1,1)$ with non-slip boundary condition. Three effects are studied in this paper: enhanced dissipation, inviscid damping and boundary layer effect. For the Navier-Stokes equations around Poiseuille flow with non-slip boundary condition, we prove that if the initial perturbation of velocity $u_{0}$ satisfies
$\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{\frac{3}{4}}$
for some constants $\epsilon_{0}>0$ independent of the viscosity $\nu$, then the solution to 2-D Navier-Stokes equations does not transit from plane Poiseuille flow. Meanwhile, for the Navier-slip boundary problem, we prove that if the initial perturbation of velocity $u_{0}$ satisfies
$\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{\frac{1}{2}}$
for some constants $\epsilon_{1}>0$ independent of the viscosity $\nu$, then the solution to 2-D Navier-Stokes equations does not transit from plane Poiseuille flow, which improves the result of [17] from $\nu^{\frac{3}{4}}$ to $\nu^{\frac{1}{2}}$.
Keywords: Poiseuille flow; transition threshold; stability.
AMS Subject Classification: 76N10, 35Q30, 35R35
1. Introduction
In this paper, we consider the 2-D incompressible Navier-Stokes equations in a channel $\Omega:=\mathbb{T}\times(-1,1)$:
$\left\{\begin{array}[]{lll}\partial_{t}v-\nu\Delta v+(v\cdot\nabla)v+\nabla q=0,\\ \nabla\cdot v=0,\\ v|_{t=0}=v_{0}(x,y),\end{array}\right.$ (1.1)
where $\nu=Re^{-1}>0$ is the viscosity, $v(t;x,y)\in\mathbb{R}^{2}$ is the velocity fields and $q\in\mathbb{R}$ is the pressure.
The plane Poiseuille flow $v_{s}=(1-y^{2},0)$ is a steady solution to the Navier-Stokes equations for constant pressure gradient $\nabla Q\equiv(q_{0},0)$. To study the stability of plane Poiseuille flow, let us introduce the perturbation $u=v-v_{s},p=q-Q$, then the perturbed equations around Poiseuille flow read as
$\left\{\begin{array}[]{lll}\partial_{t}u-\nu\Delta u+(u\cdot\nabla)u+(1-y^{2})\partial_{x}u+\left(\begin{array}[]{c}-2yu_{2}\\ 0\end{array}\right)+\nabla p=0,\\ \nabla\cdot u=0,\\ u|_{t=0}=u_{0}(x,y),\end{array}\right.$ (1.2)
$p$ is the perturbed pressure.
It is necessary to impose the boundary condition for (1.2). In this paper, we consider two types of boundary conditions:
Case I. non-slip boundary condition:
$u(t;x,\pm 1)=0;$ (1.3)
and
Case II. Navier-slip boundary condition:
$\omega(t;x,\pm 1)=0.$ (1.4)
where $\omega=\partial_{y}u_{1}-\partial_{x}u_{2}$ is the vorticity of the velocity.
The above problem (1.2) can be rewritten as
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu\Delta\omega+(1-y^{2})\partial_{x}\omega+2\partial_{x}\Phi=-\nabla\cdot(u\omega),\\ \Delta\Phi=\omega,\ u=(\partial_{y},-\partial_{x})\Phi,\\ u|_{t=0}=u_{0}(x,y),\end{array}\right.$ (1.5)
where $\Phi$ is the stream function and $\omega=\partial_{y}u_{1}-\partial_{x}u_{2}$ is the vorticity. In addition, the boundary conditions (1.3) and (1.4) are rewritten as follows respectively
$(\Phi,\partial_{y}\Phi)(t;x,\pm 1)=0$ (1.6)
and
$(\Phi,\omega)(t;x,\pm 1)=0.$ (1.7)
Our interest in this paper is to understand the long time dynamics behavior of (1.5) with (1.6), and (1.5) with (1.7) respectively for $0<\nu\ll 1$.
The hydrodynamics stability and behavior with large Reynolds number have been one of important topic in fluid dynamics since Reynolds’ experiment [37]. In the experiment, Reynolds introduced the Reynolds number $Re$ and found that the laminar flow would transit into turbulence with the increasing of the Reynolds number. The experiment suggests that the stability of the laminar flow is very sensitive to Reynolds number. Indeed, by the standard spectrum analysis, most laminar flows (e.g., plane Couette flow [38]) are stable for any Reynolds number, however the experiments indicate that the plane laminar flow would transit at high Reynolds regime [39, 48], which is so-called subcritical transition [15]. And the contradiction between the experiment and spectrum analysis is well known as Sommerfeld paradox, see [28, 39, 48] and references therein.
To understand the mechanism of transition or stability of laminar flows, there are lots of aspects to resolve the Sommerfeld paradox, including the classical works by Kelvin, Orr, Sommerfeld, C. C. Lin, see [25, 29, 34, 39, 40, 48] for instance. In particular, Kelvin [25] pointed out that the basin of attraction of laminar flows may shrink as the Reynolds number tends to infinity, which will lead nonlinear instability. With this view, Trefethen et al. [41] proposed the following transition threshold problem, whose mathematical version is formulated in [5] as follows:
Given a norm $\|\cdot\|_{X}$, determine a $\gamma=\gamma(X)$ such that
$\|u_{0}\|_{X}\leq\nu^{\gamma}\Longrightarrow\ \ stability,$
$\|u_{0}\|_{X}\gg\nu^{\gamma}\Longrightarrow\ \ instability.$
The exponent $\gamma$ is referred to as the transition threshold in the applied literatures. There are lots of works in physical or applied mathematical views devoted to determining $\gamma$ for some important shear flows, such as Couette flow and Poiseuille flow [31, 35, 36, 48]. However, the works about the rigorous analysis for transition threshold for shear flows are still needed and few.
To study the transition threshold problem, it is necessary to understand the linear stabilization effects of shear flows. In this paper, two stabilization effects of plane Poiseuille flow will be focused on: inviscid damping and enhanced dissipation. These two effects play the key roles in the hydrodynamics stability of shear flows. Inviscid damping and enhanced dissipation were firstly observed by Orr [34] and Kelvin [25] respectively, and these two effects for plane Couette flow even monotone flow are well known [2, 3, 4, 5, 6, 7, 8, 10, 12, 13, 49, 50] for different settings and domains. The results about nonlinear inviscid damping for monotone flow in Euler system can be found in [24, 33].
The inviscid damping and enhanced dissipation for non-monotone shear flow such as plane Poiseuille flow and Kolmogorov flow are more challenging. The inviscid damping and enhanced dissipation of Kolmogorov flow have been studied in [1, 21, 22, 27, 30, 46] via different methods. For the non-monotone shear flow, the linear inviscid damping and vorticity depletion for Euler system have been studied by Wei, Zhang and Zhao [45], see also Ionescu et al. [23]. We refer to [1, 20, 26] for more related results.
For the plane Poiseuille flow with high Reynolds number, the enhanced dissipation in $\mathbb{T}\times\mathbb{R}$ was firstly obtained by Coti Zelati et al. [14] via the hypocoercivity method, see also [18]. For plane Poiseuille flow in a channel $\mathbb{T}\times[-1,1]$ with Navier-slip boundary condition, Ding and Lin [17] obtained the enhanced dissipation via resolvent estimate. The linear stability for pipe Poiseuille flow has been proved by Chen, Wei and Zhang [11], in which the enhanced dissipation was also obtained in fact by the resolvent estimates.
To understand the nonlinear dynamics stability of shear flow, it is necessary to study the transition threshold problem mentioned before. There are lots of works devoted to determining the exponent $\gamma$ for different shear flows via the numerical or asymptotic analysis, see [15, 31, 36] and references therein. Recently, there are many progresses about the rigorous mathematical analysis on the transition threshold problem for different shear flows, including the Couette flow and Kolmogorov flow. For 3-D plane Couette flow, either $X$ is Gevrey class or Sobolev space, the transition threshold satisfies $\gamma\leq 1$ for the case without boundaries, see [2, 3, 4, 47]. The result $\gamma\leq 1$ also holds in a finite channel with non-slip boundary condition [12]. The transition threshold for 2-D Couette flow is proved as $\gamma\leq\frac{1}{2}$ in a channel with non-slip boundary condition [10], and $\gamma\leq\frac{1}{3}$ in $\mathbb{T}\times\mathbb{R}$ [32, 44]. More interesting related results can be found in [1, 5, 7, 20, 42].
The results for non-monotone flow are few. The transition threshold for Kolmogorov flow is proved as $\gamma\leq\frac{2}{3}+$ for 2-D domain [46] and $\gamma\leq\frac{7}{4}$ for 3-D domain [27]. For 2-D plane Poiseuille flow in $\mathbb{T}\times\mathbb{R}$, Coti Zelati et al. [14] proved that transition threshold is $\gamma\leq\frac{3}{4}+$ via hypocoercivity method, which has been improved to $\gamma\leq\frac{2}{3}+$ by Del Zotto [18]. For 2-D plane Poiseuille flow in a channel with Navier-slip boundary condition, Ding and Lin [17] proved that transition threshold is $\gamma\leq\frac{3}{4}$ by resolvent estimate method. Recently, for 3-D plane Poiseuille flow in a channel $\mathbb{T}\times[-1,1]\times\mathbb{R}$ with non-slip boundary condition, Chen, Ding, Lin and Zhang [9]proved that transition threshold is $\gamma\leq\frac{7}{4}$, which is accordant with the numerical result by Lundbladh et al. [31].
To state our main results, let us define
$P_{0}f(t,y)=\overline{f}(t,y)=\int_{\mathbb{T}}f(t,x,y)\mathrm{d}x,\ f_{\not=}=f-\overline{f}=\sum_{k\in\mathbb{Z}\setminus\{0\}}f_{k}(t,y)e^{ikx},$
where $f_{k}(t,y)$ is the Fourier transform of $f$ in $x$.
The first main result is stated as follows.
Theorem 1.1.
Suppose that $\mathrm{div}\ u_{0}=0$ and $\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{3/4}$ for some constants $\epsilon_{0}$ independent of $\nu$, then the solution to (1.5) with (1.6) is global in time with
$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\leq C\nu^{3/4},$ (1.8)
where the energy functional $\mathcal{F}_{k}$ is defined as
$\mathcal{F}_{k}:=\left\{\begin{array}[]{lll}\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}\\ \qquad\qquad\qquad+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}),&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.9)
Remark 1.1.
Let us recall that in [17], the transition threshold with $\gamma\leq\frac{3}{4}$ has been proved for the Navier-slip boundary value problem. Our Theorem 1.1 obtain the same transition threshold with $\gamma\leq\frac{3}{4}$ for the case of non-slip boundary value problem. This result implies that the boundary layer effect induced by non-slip boundary condition is not the main effect resulting in the strong instability. It is interesting to prove the stability threshold $\gamma\leq\frac{1}{2}$ mentioned in [9].
Remark 1.2.
To our knowledge, the Theorem 1.1 is the first result trying to answer the problem proposed in [9], which focuses on the transition threshold for 2-D plane Poiseuille flow with non-slip boundary condition. In our work, the stability mechanisms (including enhanced dissipation, inviscid damping) play the key role. The result $\gamma\leq\frac{3}{4}$ can be viewed as the consequence of the balance between the two good effect and boundary layer effect.
Remark 1.3.
In the energy (1.9), the weight $(1-y^{2})^{\frac{1}{2}}$ in $\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}$ is introduced to correct the singularity near boundaries $\{y=\pm 1\}$, which is due to the boundary layer effect resulted from the non-slip boundary condition. And the $|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping effect of Poiseuille flow.
The second result is to improve the transition threshold of Navier-slip boundary value problem studied in [17]. Precisely, we improve the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$, which is stated as the following theorem.
Theorem 1.2.
Suppose that $\mathrm{div}u_{0}=0,u_{0}\in H^{\frac{3}{2}+}$ with $\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{1/2}$ for some suitable small $0<\epsilon_{1}\ll 1$ independent of $\nu$, then (1.5) with (1.4) admits a global in time solution $\omega$ with
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq C\nu^{1/2},$ (1.10)
where the energy functional $\mathcal{E}_{k}$ is defined as
$\mathcal{E}_{k}:=\left\{\begin{array}[]{lll}\|\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}},&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.11)
Remark 1.4.
Indeed, by the stability estimate (1.10), we can deduce that
$\sum_{k\in\mathbb{Z}}\|u_{k}\|_{L^{\infty}L^{\infty}}\leq C\nu^{\frac{1}{2}},$
which implies that the 2-D plane Poiseuille flow is absolutely stable.
Remark 1.5.
Even for the Navier-slip boundary value problem, it is not sure that the transition threshold $\gamma\leq\frac{1}{2}$ is optimal or not. It is very interesting and challenging to verify the sharpness of $\frac{1}{2}$.
Let us give some comments about our results.
To our knowledge, our Theorem 1.1 firstly gives the transition threshold for the 2-D plane Poiseuille flow with non-slip boundary condition, i.e., $\gamma\leq\frac{3}{4}$. Compared with [17], our result implies that the boundary layer effect induced by non-slip boundary condition is not the main mechanism resulting in the strong nonlinear instability. The linear effects are formulated in the functional (1.9), including the enhanced dissipation, inviscid damping and boundary layer effect. The key point is the space-time estimates for linearized Navier-Stokes system with source terms, which corresponding to the careful resolvent estimates of Orr-Sommerfeld equation around plane Poiseuille flow established in [9].
In addition, compared with [17], the new trouble in this paper is boundary layer effect. Note that near boundaries, it holds that $1-y^{2}\sim 1-|y|$. To control the boundary layer effect, the weight estimate with weight $1-y^{2}$ is needed, which avoid the loss of $\nu$ in $L^{\infty}L^{2}$-estimate for vorticity. If one only considers the $L^{\infty}L^{2}$-estimate for vorticity without weight $1-y^{2}$, there is loss of $\nu^{\frac{1}{4}}$ in the estimate, which is resulted from the boundary layer effect. In addition, to achieve the transition threshold $\gamma\leq\frac{3}{4}$ as in [17], the enhanced dissipation and inviscid damping effect are needed, which contribute the stabilizing mechanism to balance the boundary layer effect.
Our second result improves the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$. Compared with [17], the main contribution to the improvement is the inviscid damping effect. Indeed, the result $\nu^{3/4}$ in [17] is due to the enhanced dissipation effect. To improve it to $\nu^{1/2}$, the key point is the inviscid damping effect and the good structure of the Rayleigh part (or inviscid part)
$(1-y^{2})+2\Delta^{-1},$
or in the Fourier version, namely,
$(1-y^{2})+2(\partial_{y}^{2}-k^{2})^{-1}.$
The above Rayleigh structure gives good positive estimates for velocity and vorticity(with weight), which along with the diffusion yield better space-time $L^{2}L^{2}$ estimate to vorticity, see Section 2 for instance. This implies the good stability mechanism of Poiseuille flow in some sense.
Moreover, with the good positive structure, one may notice that our proof is different from that of [17] and [14, 18]. The proof for Theorem 1.2 relies on a direct energy argument with some positive structure for the toy model around Poiseuille flow, see Section 2 for formal analysis.
Let us give some notations to end this section. Throughout this paper, we always assume that $|k|\geq 1$ and denote by $C$ a positive constant independent of $\nu,k$. We also use $a\sim b$ for $C^{-1}b\leq a\leq Cb$ and $a\lesssim b$ for $a\leq Cb$ for some constants $C>0$ independent of $\nu,k$.
2. Sketch and key ideas of the proof
In this section, let us give the sketch and some key ideas in our analysis.
2.1. The stability for Poiseuille flow with non-slip boundary condition
To obtain the transition threshold of plane Poiseuille flow with non-slip boundary condition, the key point is to establish the space-time estimates for the following linearized system
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\partial_{y}\varphi(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$
where $f_{i}\in L^{2}$ and $k\not=0$. The basic space-time estimates are the $L^{2}L^{2}$ estimates for vorticity and velocity field, which are due to enhanced dissipation and inviscid damping induced by plane Poiseuille flow. See Section 3 and [9] for details.
Compared with the Navier-slip boundary condition in [17], the new effect may result in instability is the boundary layer effect, which is resulted from the non-slip boundary condition. Precisely, due to the singularity occurring near the boundaries, there is loss of $\nu$ in the space-time estimates if following the arguments of [17], which will lead to a threshold worse than $\nu^{3/4}$. To overcome the difficulty, the main idea is to introduce the suitable weight to control this effect near boundaries $\{y=\pm 1\}$. Since near $\{y=\pm 1\}$, it holds that $1-y^{2}\sim 1-|y|$, the suitable weight in our argument is chosen as $1-y^{2}$. Precisely, we deduce the weighed estimate as follows:
$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2},$
see Section 3 for details. There are some advantages about this weight: it corrects the behavior of the vorticity near boundaries and it can be viewed as the good weight in some Hardy-type inequalities. If we consider the estimates without the weight, we only obtain that
$\nu^{\frac{1}{2}}\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu^{\frac{3}{2}}\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2},$
There is loss of $\nu$ in the above estimates, and the weight $1-y^{2}$ avoid this loss of $\nu$.
With the above arguments, the space-time estimates for nonlinear stability read as
$\mathcal{F}_{k}\lesssim\|\omega_{0}\|_{H^{2+}}+\mathrm{nonlinear\ terms},$
where the energy functional for $k\not=0$ is defined as
$\mathcal{F}_{k}:=\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}).$
The term $\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}$ is due to boundary layer effect, $(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}$ is due to enhanced dissipation, and $|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}}$ is due to inviscid damping. In other words, the result is the consequence of the balance between the enhanced dissipation, inviscid damping and boundary layer effect.
Then the nonlinear stability follows from continuous argument, see Subsection 5.1 for details.
2.2. The improved transition threshold with Navier-slip boundary condition
In our analysis, compared with [17], the most important point is the improved space-time estimates. Indeed, in the view of Fourier transform, the toy model for our problem (linear and nonlinear sense) reads
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ u=(\partial_{y},-ik)\varphi,\\ (\varphi,\omega)|_{y=\pm 1}=(0,0),\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (2.1)
where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$.
Let us recall that the space-time estimates in [17] (in the Fourier version) read as
$\|e^{c^{\prime}(\nu|k|)^{1/2}t}\omega\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{1/2}$ $\|e^{c^{\prime}(\nu|k|)^{1/2}t}\omega\|_{L^{2}L^{2}}^{2}+\nu\|e^{c^{\prime}(\nu|k|)^{1/2}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c^{\prime}(\nu|k|)^{1/2}t}(f_{1},f_{2})\|_{L^{2}L^{2}}^{2},$
where the term $(\nu|k|)^{1/2}\|e^{c^{\prime}(\nu|k|)^{1/2}t}\omega\|_{L^{2}L^{2}}^{2}$ is due to enhanced dissipation, which is resulted from the resolvent estimate obtained in [17]. In a word, the stabilization mechanism used in [17] is only dissipation, including the enhanced dissipation and heat diffusion.
To establish the improved space-time estimates, we consider the energy form as follows
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}$ $\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}$ (2.2)
$+ik\langle(1-y^{2})\omega+2\varphi,\omega\rangle=\langle-ikf_{1}-\partial_{y}f_{2},\omega\rangle:=\mathrm{RHS}.$
The first key observation is the inviscid Rayleigh part (imaginary part) of (2.2) will give the space-time estimate for velocity field, which corresponds to the inviscid damping effect (integral form) of velocity field. More precisely, the Rayleigh part yields the following positive estimate (see Lemma 4.2)
$\|(\varphi^{\prime},|k|\varphi)\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\lesssim\langle(1-y^{2})w+2\varphi,w\rangle,$
which implies the following improved estimate
$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\left(\|\varphi^{\prime}\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\right)$ (2.3)
$\lesssim$ $\mathrm{Re}\langle(1-y^{2})\omega+2\varphi,\omega\rangle\lesssim|k|^{-1}|\mathrm{Im}\ \mathrm{RHS}|,$
which implies the estimate for velocity field. See the proof of Proposition 4.2 for details.
It should be pointed out that the estimate (2.3) is resulted from the Rayleigh part and it is valid uniform in $\nu$, and this implies the intrinsic structure and stabilization mechanism of the Poiseuille flow.
The second key observation is the improvement of the $L^{2}L^{2}$-estimate for $\omega$. Precisely, with (2.3) at hands, by applying Lemma 4.1, we have
$(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}\lesssim(\nu|k|^{2})^{1/3}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{4/3}\|\partial_{y}\omega\|_{L^{2}}^{2/3}\lesssim\nu\|\partial_{y}\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}$
for $\omega|_{y=\pm 1}=0$. Then we establish
$\frac{\mathrm{d}}{\mathrm{d}t}$ $\|\omega\|_{L^{2}}^{2}+c(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}+\cdots\lesssim|\mathrm{RHS}|,$ (2.4)
which suggests that the $L^{2}L^{2}$-estimate for $\omega$ in [17] can be improved as
$\|e^{c_{0}(\nu|k|^{2})^{1/3}t}\omega\|_{L^{\infty}L^{2}}^{2}+(\nu|k|^{2})^{1/3}\|e^{c_{0}(\nu|k|^{2})^{1/3}t}$ $\omega\|_{L^{2}L^{2}}^{2}+\cdots$ (2.5)
$\lesssim\|\omega_{0}\|_{L^{2}}^{2}+\mathrm{source\ terms}.$
Therefore, the improved space-time estimates for toy model (2.1) follow. See Proposition 4.2 for details.
The nonlinear stability and transition threshold are also based on the improved space-time estimates in Proposition 4.2. Indeed, in [17], we only used the enhanced dissipation effect to obtain that transition threshold $\nu^{\gamma}$ with $\gamma\leq 3/4$. In this paper, additional to the enhanced dissipation, the inviscid damping for velocity field is established. With the effects and Proposition 4.2, the space-time estimates of the $x$-nonzero modes (i.e., in the Fourier version with $k\not=0$) for nonlinear problem read as
$\mathcal{E}_{k}\lesssim\|\omega_{0,k}\|_{L^{2}}+\mathrm{nonlinear\ terms},$
where the energy functional $\mathcal{E}_{k}$ with $k\not=0$ is defined as
$\mathcal{E}_{k}=\|\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}}.$
Compared with [17], the term $(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}$ is due to the improved space-time estimates, see the formal analysis (2.4)-(2.5). If following the arguments of [17], we can only obtain the bound of $(\nu|k|^{2})^{1/4}\|\omega_{k}\|_{L^{2}L^{2}}$, which is worse than the bound obtained in this paper. In addition, the term $|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping. It should be noted that it is independent of $\nu$ and it would avoid the loss of $\nu$ in nonlinear interaction. See Subsection 5.2 for details.
With the above key points, one can improve the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$.
3. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow: non-slip boundary condition
The linearized Navier-Stokes equations around Poiseuille flow read as
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\partial_{y}\varphi(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (3.1)
where $f_{i}\in L^{2}$ and $k\not=0$.
The space-time estimates for (3.1) are stated as follows.
Proposition 3.1.
Suppose the $\omega_{0}\in H^{1}_{0}$ with $\langle\omega_{0},e^{\pm|k|y}\rangle=0$, then the solution $\omega$ to (3.1) enjoys that the following estimates:
(1) If $\nu|k|^{2}\leq 2$, it holds that
$(\nu|k|)^{\frac{1}{2}}$ $\|\omega\|_{L^{2}L^{2}}^{2}+|k|(\|u\|_{L^{\infty}L^{2}}^{2}+\|u\|_{L^{2}L^{2}}^{2})+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}L^{2}}^{2}$ (3.2)
$+\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
(2) If $\nu|k|^{2}\geq 2$, it holds that
$\|\omega\|_{L^{\infty}L^{2}}^{2}$ $+|k|^{2}(\|u\|_{L^{\infty}L^{2}}^{2}+\|u\|_{L^{2}L^{2}}^{2})+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$ (3.3)
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
Proof.
Indeed, it remains to give the space-time estimates of $\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}$ and $|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$ for $\nu|k|^{2}\leq 2$, other estimates can be found in [9].
By testing (3.1) by $(1-y^{2})\omega$, we have
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+\nu\int_{-1}^{1}(-2y)\partial_{y}\omega\overline{\omega}\mathrm{d}y$
$+ik\|(1-y^{2})\omega\|_{L^{2}}^{2}=\langle-ikf_{1},\omega\rangle+\langle f_{2},\partial_{y}((1-y^{2})\omega)\rangle+2ik\langle\varphi,(1-y^{2})\omega\rangle,$
which gives that
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}$
$\leq$ $2\nu\|\omega\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}+2|k|\|\varphi\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}$
$+|k|\|f_{1}\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}((1-y^{2})\omega)\|_{L^{2}}$
$\leq$ $\nu^{\frac{1}{2}}\|\omega\|_{L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}+4|k|\|\varphi\|_{L^{2}}^{2}$
$+\frac{1}{2}\nu|k|^{2}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\frac{1}{2}(\nu|k|^{2})^{-1}\|(1-y^{2})^{\frac{1}{2}}f_{1}\|_{L^{2}}^{2}$
$+\frac{1}{2}\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{2}\nu\|(1-y^{2})^{\frac{1}{2}}\partial_{y}\omega\|_{L^{2}}^{2}$
or
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}$
$\leq$ $2\nu\|\omega\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}+2|k|\|\varphi\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}$
$+|k|\|f_{1}\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}((1-y^{2})\omega)\|_{L^{2}}$
$\leq$ $\nu^{\frac{1}{2}}\|\omega\|_{L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}+4|k|\|\varphi\|_{L^{2}}^{2}$
$+\frac{1}{2}|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})^{\frac{1}{2}}f_{1}\|_{L^{2}}^{2}$
$+\frac{1}{2}\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{2}\nu\|(1-y^{2})^{\frac{1}{2}}\partial_{y}\omega\|_{L^{2}}^{2}.$
Therefore, we obtain that
$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\nu^{\frac{1}{2}}\|\omega\|_{L^{2}L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|u\|_{L^{2}L^{2}}^{2}$
$+\min\{(\nu|k|^{2})^{-1},|k|\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
The proof is completed. ∎
4. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow: Navier-slip boundary condition
In this section, we will establish the space-time estimates for linearized Navier-Stokes equations as follows
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\omega(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (4.1)
where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$. It is noted that $\omega_{0}(k,\pm 1)=0$ ensures $\omega(t,k,\pm 1)=0$ for any $t\geq 0$.
Before giving the main result of this section, let us recall two lemmas from [9].
Lemma 4.1.
For $f\in H^{1}_{0}(-1,1)$, it holds that
$\|f\|_{L^{2}}^{2}\leq\|(1-y^{2})^{\frac{1}{2}}f\|_{L^{2}}^{\frac{4}{3}}\|(1-y^{2})^{-1}f\|_{L^{2}}^{\frac{2}{3}}\leq C\|(1-y^{2})^{\frac{1}{2}}f\|_{L^{2}}^{\frac{4}{3}}\|\partial_{y}f\|_{L^{2}}^{\frac{2}{3}}.$
Lemma 4.2.
Suppose that $(\partial_{y}^{2}-|k|^{2})\varphi=w,\varphi(\pm 1)=0,|k|\neq 0$. Then it holds that
$2\left(\bigg{\|}\varphi^{\prime}+\dfrac{2y\varphi}{1-y^{2}}\bigg{\|}_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\right)\leq\langle(1-y^{2})w+2\varphi,w\rangle,$
$\|(\varphi^{\prime},|k|\varphi)\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\leq\big{(}8/|k|^{2}+1\big{)}\langle(1-y^{2})w+2\varphi,w\rangle.$
The space-time estimates for (4.1) are stated as follows.
Proposition 4.1.
Let $\omega$ be the solution to (4.1) with $\omega_{0}\in H^{1}_{0}(-1,1)$, and $f_{i}\in L^{2}(i=1,2)$, then there exists $0<c_{0}\ll 1$ such that
$\|\omega\|_{L^{\infty}L^{2}}^{2}+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$ (4.2)
$\qquad+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}$
and
$\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{\infty}L^{2}}^{2}$ (4.3)
$\qquad+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}L^{2}}^{2}$
$\qquad+|k|\left(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}L^{2}}^{2}.$
Proof.
The standard energy arguments give that
$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+ik\langle(1-y^{2})\omega+2\varphi,\omega\rangle$
$=$ $\langle-ikf_{1},\omega\rangle+\langle f_{2},\partial_{y}\omega\rangle=:\mathrm{RHS}.$
By Lemma 4.2, we have
$|k|$ $\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}\lesssim|\mathrm{Im}\ \mathrm{RHS}|,$
which gives that
$|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}\lesssim|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}+|\mathrm{Im}\ \mathrm{RHS}|\lesssim|\mathrm{Im}\ \mathrm{RHS}|.$
Therefore, we have
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\frac{\nu}{2}\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}}^{2}$
$\qquad\qquad+\frac{|k|}{2}(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
$\leq$ $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}$
$\qquad+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}\right)\leq C|\mathrm{RHS}|,$
where we have used Lemma 4.1 for $\omega(\pm 1)=0$.
The RHS is bounded as
$|\mathrm{RHS}|\leq$ $|k|\|f_{1}\|_{L^{2}}\|\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}$
$\leq$ $\frac{1}{4}\left(\nu|k|^{2}\|\omega\|_{L^{2}}^{2}+\nu\|\partial_{y}\omega\|_{L^{2}}^{2}\right)+C\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+\frac{1}{4}|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2},$
or
$|\mathrm{RHS}|\leq$ $|k|\|f_{1}\|_{L^{2}}\|\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}$
$\leq$ $\epsilon_{1}(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}+C(\nu|k|^{2})^{-1/3}|k|^{2}\|f_{1}\|_{L^{2}}^{2}$
$+\frac{1}{4}\nu\|\partial_{y}\omega\|_{L^{2}}^{2}+C\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{4}|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}.$
Taking $\epsilon_{1}\ll\min\{c,1\}$, we have
$\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}}^{2}$
$\qquad\qquad\qquad+|k|(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
$\lesssim$ $\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}}^{2}.$
Therefore, we get that
$\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}L^{2}}^{2}$
$\qquad\qquad\qquad+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}.$
Moreover, we have
$\frac{\mathrm{d}}{\mathrm{d}t}$ $\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}$
$+|k|(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
$\lesssim$ $\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}}^{2}$
$+c_{0}(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}.$
Taking $c_{0}\ll 1$, we have
$\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}L^{2}}^{2}$
$+|k|\left(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}L^{2}}^{2}.$
The proof is completed. ∎
5. Nonlinear stability
In this section, we will give the proof of the main results, i.e., Theorem 1.1 and Theorem 1.2.
5.1. Nonlinear stability with non-slip boundary condition: Proof of Theorem 1.1
Recall that the Navier-Stokes equations around Poiseuille flow with non-slip boundary condition read as
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}-\nu(\partial_{y}^{2}-k^{2})\omega_{k}+ik(1-y^{2})\omega_{k}+2ik\varphi_{k}=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\ \ u_{k}=(\partial_{y},-ik)\varphi_{k},\\ \varphi_{k}(t,k,\pm 1)=\partial_{y}\varphi_{k}(t,k,\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.1)
where
$(f_{1})_{k}=-\sum_{l\in\mathbb{Z}}(u_{1})_{l}\omega_{k-l},\ (f_{2})_{k}=-\sum_{l\in\mathbb{Z}}(u_{2})_{l}\omega_{k-l}.$
Let us recall that
$\mathcal{F}_{k}:=\left\{\begin{array}[]{lll}\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}\\ \qquad\qquad\qquad+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}),&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$
Now we are ready to prove Theorem 1.1.
Proof of Theorem 1.1.
The global well-posedness of 2-D Navier-Stokes equation around the Poiseuille flow can be obtained by classical theory for Navier-Stokes equation. Here we only focus on the stability estimates. It should be noted that the case of $\nu|k|^{2}\geq 2$ is trivial by using (3.3) and we only consider the case of $\nu|k|^{2}\leq 2$.
Let us first consider the zero modes. If $k=0$, by the divergence-free condition, we have $\partial_{y}(u_{2})_{0}=0$, which implies by $(u_{2})_{0}|_{y=\pm 1}=0$ that $(u_{2})_{0}=0$. Therefore, the equation for zero modes is reduced to
$\partial_{t}(u_{1})_{0}-\nu\partial_{y}^{2}(u_{1})_{0}=-\partial_{y}(f_{2})_{0}.$
By the similar arguments as for 2-D Couette flow [10], we have
$\mathcal{F}_{0}^{2}\lesssim\nu^{-1}\|(f_{2})_{0}\|_{L^{2}L^{2}}^{2}+\|\overline{\omega_{0}}\|_{L^{2}}^{2}.$ (5.2)
For $\mathcal{F}_{k}$, by Proposition 3.1, one has
$\mathcal{F}_{k}\lesssim|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}.$
For $k\not=0$, it holds that
$\|(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$ (5.3)
$\leq$ $\sum_{l\in\mathbb{Z}}\left\|(u_{2})_{l}\right\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}$
$\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\|\partial_{y}(u_{2})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
$\lesssim$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}(|l|\|(u_{1})_{l}\|_{L^{\infty}L^{2}})^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
$\lesssim$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}(|l|^{\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
$\lesssim$ $\nu^{-\frac{1}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}$
and
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.4)
$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|(u_{1})_{l}\omega_{k-l}\|_{L^{2}L^{2}}.$
It remains to estimate the terms on the right-hand side of (5.9).
For $y\in[-1,0]$, it holds that $(1-|y|)=(1+y)\leq(1+y)(1-y)=1-y^{2}$; and for $y\in[0,1]$, $(1-|y|)=(1-y)\leq(1+y)(1-y)=1-y^{2}$, thus $1-y^{2}\geq 1-|y|$. For $l\not=0,k$, it holds that $|k|\lesssim|l||k-l|$, then
$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|(u_{1})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\left\|\frac{(u_{1})_{l}}{(1-y^{2})^{\frac{1}{2}}}\right\|_{L^{2}L^{\infty}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|\partial_{y}(u_{1})_{l}\|_{L^{2}L^{2}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|\omega_{l}\|_{L^{2}L^{2}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}(\nu|l|)^{-\frac{1}{4}}\mathcal{F}_{l}\mathcal{F}_{k-l}$
and
$\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}$ $\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}\|\omega_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{4}}|k|^{-\frac{1}{4}}\mathcal{F}_{0}\mathcal{F}_{k}.$
Thus, we have
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{4}}|k|^{-\frac{1}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}$ (5.5)
Combining (5.2), (5.3) and (5.5), we have
$\mathcal{F}_{k}\lesssim$ $|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}$
$\lesssim$ $|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l},$
$\mathcal{F}_{0}\lesssim$ $\|\overline{\omega}_{0}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}\setminus\{0\}}\mathcal{F}_{l}\mathcal{F}_{-l},$
which gives that
$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\lesssim\sum_{k\in\mathbb{Z}}|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{4}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}.$
Due to $\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{\frac{3}{4}}$, then it holds that
$\sum_{k\in\mathbb{Z}}|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}\leq\epsilon_{0}C\nu^{\frac{3}{4}}.$
If $\epsilon_{0}$ is suitably small, one can deduce from continuous argument that
$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\leq\epsilon_{0}C\nu^{\frac{3}{4}}.$
The proof is completed. ∎
5.2. Nonlinear stability with Navier-slip boundary condition: Proof of Theorem 1.2
For the nonlinear problem with Navier-slip boundary condition, let us recall that
$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}-\nu(\partial_{y}^{2}-k^{2})\omega_{k}+ik(1-y^{2})\omega_{k}+2ik\varphi_{k}=-ik(f_{1})_{k}-\partial_{y}(f_{2})_{k},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\ \varphi_{k}(\pm 1)=\omega_{k}(\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.6)
where
$(f_{1})_{k}=-\sum_{l\in\mathbb{Z}}(u_{1})_{l}\omega_{k-l},\ (f_{2})_{k}=-\sum_{l\in\mathbb{Z}}(u_{2})_{l}\omega_{k-l}.$
For simplicity, we define $\overline{\omega_{0}}:=\omega_{0,0}(0,y)$ for the zero mode part of the initial data.
Recall that for $k\not=0,$
$\mathcal{E}_{k}=(\nu|k|^{2})^{\frac{1}{6}}\|\omega_{k}\|_{L^{2}L^{2}}+\|\omega_{k}\|_{L^{\infty}L^{2}}+|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}},$
and $\mathcal{E}_{0}=\|\overline{\omega}\|_{L^{\infty}L^{2}}$ for $k=0$.
Now we are on a position to give the proof of Theorem 1.2.
Proof of Theorem 1.2.
The global well-posedness of 2-D Navier-Stokes equation around the Poiseuille flow can be obtained by classical theory for Navier-Stokes equation. So we only focus on the stability estimates in Theorem 1.2.
First let us treat with the zero modes. If $k=0$, by the divergence-free condition, we have $\partial_{y}(u_{2})_{0}=0$, which implies by $(u_{2})_{0}|_{y=\pm 1}=0$ that $(u_{2})_{0}=0$. Therefore, the equation for zero modes is reduced to
$\partial_{t}(u_{1})_{0}-\nu\partial_{y}^{2}(u_{1})_{0}=-\partial_{y}(f_{2})_{0}.$
By the similar arguments as for 2-D Couette flow [10], we have
$\mathcal{E}_{0}^{2}\lesssim\nu^{-1}\|(f_{2})_{0}\|_{L^{2}L^{2}}^{2}+\|\overline{\omega_{0}}\|_{L^{2}}^{2}.$ (5.7)
For $\mathcal{E}_{k}$, by the space-time estimates in Proposition 4.2, one has
$\mathcal{E}_{k}\lesssim\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{3}}|k|^{\frac{1}{3}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}},$
where we have used the fact that $\min\{(\nu^{-1}|k|^{4})^{\frac{1}{3}},\nu^{-1}\}\leq(\nu|k|^{-1})^{-\frac{2}{3}}$.
For $k\not=0$, it holds that
$\|(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}\leq\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{2}L^{\infty}}\|\omega_{k-l}\|_{L^{\infty}L^{2}}$ (5.8)
$\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|\partial_{y}(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\mathcal{E}_{k-l}$
$\leq$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{E}_{l})^{\frac{1}{2}}(|l|\|(u_{1})_{l}\|_{L^{2}L^{2}})^{\frac{1}{2}}\mathcal{E}_{k-l}$
$\leq$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{E}_{l})^{\frac{1}{2}}(|l|^{-\frac{1}{2}}|l|\mathcal{E}_{l})^{\frac{1}{2}}\mathcal{E}_{k-l}\lesssim\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}$
and
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.9)
$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}.$
It remains to estimate the terms on the right-hand side of (5.9).
For $l\not=0,k$, it holds that $|k|\lesssim|l||k-l|$, then
$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}$
$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}|l|^{-\frac{1}{2}}\mathcal{E}_{l}\nu^{-\frac{1}{6}}|k-l|^{-\frac{1}{3}}\mathcal{E}_{k-l}$
$\lesssim$ $\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\mathcal{E}_{l}\mathcal{E}_{k-l}$
and
$\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}$ $\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}\|\omega_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\mathcal{E}_{0}\mathcal{E}_{k}.$
Thus, we have
$\|(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}$ (5.10)
Combining (5.7), (5.8) and (5.10), we have
$\mathcal{E}_{k}\lesssim$ $\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{3}}|k|^{\frac{1}{3}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}$
$\lesssim$ $\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l},$
$\mathcal{E}_{0}\lesssim$ $\|\overline{\omega}_{0}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}\setminus\{0\}}\mathcal{E}_{l}\mathcal{E}_{-l},$
which gives that
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\lesssim\sum_{k\in\mathbb{Z}}\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}.$
Due to $\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{1/2}$, then one has
$\sum_{k\in\mathbb{Z}}\|\omega_{0,k}\|_{L^{2}}\leq\epsilon_{1}C\nu^{\frac{1}{2}}.$
If $\epsilon_{1}$ is suitably small, one can deduce from continuous argument that
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq\epsilon_{1}C\nu^{\frac{1}{2}}.$
The proof is completed. ∎
Acknowledgment.
The authors thank to Prof. Zhifei Zhang for valuable discussions and suggestions. S. Ding is supported by the Key Project of National Natural Science Foundation of China under Grant 12131010, National Natural Science Foundation of China under Grant 12271032 and Natural Science Foundation of Guangdong Province under Grant 2021A1515010249 and 2021A1515010303. Z. Lin is partially supported by National Natural Science Foundation of China under Grant 11971009.

### 03 Mar 2024 v2

@@
 Abstract.
-In this paper, we study the stability for 2-D plane Poiseuille flow $(1-y^{2},0)$ in a channel $\mathbb{T}\times(-1,1)$ with non-slip boundary condition.
-Three effects are studied in this paper: enhanced dissipation, inviscid damping and boundary layer effect.
-For the Navier-Stokes equations around Poiseuille flow with non-slip boundary condition, we prove that if the initial perturbation of velocity $u_{0}$ satisfies
-$\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{\frac{3}{4}}$
-for some constants $\epsilon_{0}>0$ independent of the viscosity $\nu$, then the solution to 2-D Navier-Stokes equations does not transit from plane Poiseuille flow.
-Meanwhile, for the Navier-slip boundary problem, we prove that if the initial perturbation of velocity $u_{0}$ satisfies
-$\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{\frac{1}{2}}$
-for some constants $\epsilon_{1}>0$ independent of the viscosity $\nu$, then the solution to 2-D Navier-Stokes equations does not transit from plane Poiseuille flow, which improves the result of [17] from $\nu^{\frac{3}{4}}$ to $\nu^{\frac{1}{2}}$.
+In this paper, we study the stability for 2-D plane Poiseuille flow $(1-y^{2},0)$ in a channel $\mathbb{T}\times(-1,1)$ with Navier-slip boundary condition.
+We prove that if the initial perturbation for velocity field $u_{0}$ satisfies that $\|u_{0}\|_{H^{\frac{7}{2}+}}\leq\epsilon_{1}\nu^{2/3}$ for some suitable small $0<\epsilon_{1}\ll 1$ independent of viscosity coefficient $\nu$, then the solution to the Navier-Stokes equations is global in time and does not transit from the plane Poiseuille flow.
+This result improves the result of [18] from $3/4$ to $2/3$.
 Keywords: Poiseuille flow; transition threshold; stability.
@@
 1. Introduction
-In this paper, we consider the 2-D incompressible Navier-Stokes equations in a channel $\Omega:=\mathbb{T}\times(-1,1)$:
+In this paper, we consider the 2-D incompressible Navier-Stokes (N-S) equations in a channel $\Omega:=\mathbb{T}\times(-1,1)$:
 $\left\{\begin{array}[]{lll}\partial_{t}v-\nu\Delta v+(v\cdot\nabla)v+\nabla q=0,\\ \nabla\cdot v=0,\\ v|_{t=0}=v_{0}(x,y),\end{array}\right.$ (1.1)
@@
 $p$ is the perturbed pressure.
-It is necessary to impose the boundary condition for (1.2). In this paper, we consider two types of boundary conditions:
-Case I. non-slip boundary condition:
-$u(t;x,\pm 1)=0;$ (1.3)
-and
-Case II. Navier-slip boundary condition:
+It is necessary to impose the boundary condition for (1.2). In this paper, we consider the Navier-slip boundary condition
 $\omega(t;x,\pm 1)=0.$ (1.3)
@@
 $\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu\Delta\omega+(1-y^{2})\partial_{x}\omega+2\partial_{x}\Phi=-\nabla\cdot(u\omega),\\ \Delta\Phi=\omega,\ u=(\partial_{y},-\partial_{x})\Phi,\\ u|_{t=0}=u_{0}(x,y),\end{array}\right.$ (1.4)
-where $\Phi$ is the stream function and $\omega=\partial_{y}u_{1}-\partial_{x}u_{2}$ is the vorticity. In addition, the boundary conditions (1.3) and (1.4) are rewritten as follows respectively
-$(\Phi,\partial_{y}\Phi)(t;x,\pm 1)=0$ (1.6)
-and
+where $\Phi$ is the stream function and $\omega=\partial_{y}u_{1}-\partial_{x}u_{2}$ is the vorticity. In addition, the boundary condition (1.3) is rewritten as follows
 $(\Phi,\omega)(t;x,\pm 1)=0.$ (1.5)
-Our interest in this paper is to understand the long time dynamics behavior of (1.5) with (1.6), and (1.5) with (1.7) respectively for $0<\nu\ll 1$.
+Our interest in this paper is to understand the long time dynamics behavior of (1.4) with (1.5) for $0<\nu\ll 1$.
 The hydrodynamics stability and behavior with large Reynolds number have been one of important topic in fluid dynamics since Reynolds’ experiment [38].
@@
 The transition threshold for 2-D Couette flow is proved as $\gamma\leq\frac{1}{2}$ in a channel with non-slip boundary condition [10], and $\gamma\leq\frac{1}{3}$ in $\mathbb{T}\times\mathbb{R}$ [33, 45].
+Recently, Coti Zelati et al.
+showed that the transition threshold for Couette flow in 3-D Bounssinesq equations enjoys $\gamma<1$ [15].
 More interesting related results can be found in [1, 5, 7, 21, 43].
@@
 where $f_{k}(t,y)$ is the Fourier transform of $f$ in $x$.
-The first main result is stated as follows.
+The main result is stated as follows, which improves the transition threshold of Navier-slip boundary value problem studied in [18].
+Precisely, we improve the result of [18] from $\nu^{3/4}$ to $\nu^{2/3}$, which is stated as the following theorem.
 Theorem 1.1.
-Suppose that $\mathrm{div}\ u_{0}=0$ and $\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{3/4}$ for some constants $\epsilon_{0}$ independent of $\nu$, then the solution to (1.5) with (1.6) is global in time with
-$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\leq C\nu^{3/4},$ (1.8)
-where the energy functional $\mathcal{F}_{k}$ is defined as
-$\mathcal{F}_{k}:=\left\{\begin{array}[]{lll}\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}\\ \qquad\qquad\qquad+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}),&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.9)
-Remark 1.1.
-Let us recall that in [17], the transition threshold with $\gamma\leq\frac{3}{4}$ has been proved for the Navier-slip boundary value problem.
-Our Theorem 1.1 obtain the same transition threshold with $\gamma\leq\frac{3}{4}$ for the case of non-slip boundary value problem.
-This result implies that the boundary layer effect induced by non-slip boundary condition is not the main effect resulting in the strong instability.
-It is interesting to prove the stability threshold $\gamma\leq\frac{1}{2}$ mentioned in [9].
-Remark 1.2.
-To our knowledge, the Theorem 1.1 is the first result trying to answer the problem proposed in [9], which focuses on the transition threshold for 2-D plane Poiseuille flow with non-slip boundary condition.
-In our work, the stability mechanisms (including enhanced dissipation, inviscid damping) play the key role.
-The result $\gamma\leq\frac{3}{4}$ can be viewed as the consequence of the balance between the two good effect and boundary layer effect.
-Remark 1.3.
-In the energy (1.9), the weight $(1-y^{2})^{\frac{1}{2}}$ in $\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}$ is introduced to correct the singularity near boundaries $\{y=\pm 1\}$, which is due to the boundary layer effect resulted from the non-slip boundary condition.
-And the $|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping effect of Poiseuille flow.
-The second result is to improve the transition threshold of Navier-slip boundary value problem studied in [17].
-Precisely, we improve the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$, which is stated as the following theorem.
-Theorem 1.2.
-Suppose that $\mathrm{div}u_{0}=0,u_{0}\in H^{\frac{3}{2}+}$ with $\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{1/2}$ for some suitable small $0<\epsilon_{1}\ll 1$ independent of $\nu$, then (1.5) with (1.4) admits a global in time solution $\omega$ with
-$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq C\nu^{1/2},$ (1.10)
+Suppose that $\mathrm{div}u_{0}=0,u_{0}\in H^{\frac{7}{2}+}$ with $\|u_{0}\|_{H^{\frac{7}{2}+}}\leq\epsilon_{1}\nu^{2/3}$ for some suitable small $0<\epsilon_{1}\ll 1$ independent of $\nu$, then (1.4) with (1.3) admits a global in time solution $\omega$ with
+$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq C\nu^{2/3},$ (1.6)
 where the energy functional $\mathcal{E}_{k}$ is defined as
-$\mathcal{E}_{k}:=\left\{\begin{array}[]{lll}\|\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}},&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.11)
+$\mathcal{E}_{k}:=\left\{\begin{array}[]{lll}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}+\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}\\ \qquad+(\nu|k|)^{1/4}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}},&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.7)
 Remark 1.1.
 Indeed, by the stability estimate (1.6), we can deduce that
-$\sum_{k\in\mathbb{Z}}\|u_{k}\|_{L^{\infty}L^{\infty}}\leq C\nu^{\frac{1}{2}},$
+$\sum_{k\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{\infty}L^{\infty}}\leq C\nu^{2/3},$
 which implies that the 2-D plane Poiseuille flow is absolutely stable.
 Remark 1.2.
-Even for the Navier-slip boundary value problem, it is not sure that the transition threshold $\gamma\leq\frac{1}{2}$ is optimal or not.
-It is very interesting and challenging to verify the sharpness of $\frac{1}{2}$.
-Let us give some comments about our results.
-To our knowledge, our Theorem 1.1 firstly gives the transition threshold for the 2-D plane Poiseuille flow with non-slip boundary condition, i.e., $\gamma\leq\frac{3}{4}$.
-Compared with [17], our result implies that the boundary layer effect induced by non-slip boundary condition is not the main mechanism resulting in the strong nonlinear instability.
-The linear effects are formulated in the functional (1.9), including the enhanced dissipation, inviscid damping and boundary layer effect.
-The key point is the space-time estimates for linearized Navier-Stokes system with source terms, which corresponding to the careful resolvent estimates of Orr-Sommerfeld equation around plane Poiseuille flow established in [9].
-In addition, compared with [17], the new trouble in this paper is boundary layer effect.
-Note that near boundaries, it holds that $1-y^{2}\sim 1-|y|$.
-To control the boundary layer effect, the weight estimate with weight $1-y^{2}$ is needed, which avoid the loss of $\nu$ in $L^{\infty}L^{2}$-estimate for vorticity.
-If one only considers the $L^{\infty}L^{2}$-estimate for vorticity without weight $1-y^{2}$, there is loss of $\nu^{\frac{1}{4}}$ in the estimate, which is resulted from the boundary layer effect.
-In addition, to achieve the transition threshold $\gamma\leq\frac{3}{4}$ as in [17], the enhanced dissipation and inviscid damping effect are needed, which contribute the stabilizing mechanism to balance the boundary layer effect.
-Our second result improves the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$.
-Compared with [17], the main contribution to the improvement is the inviscid damping effect.
-Indeed, the result $\nu^{3/4}$ in [17] is due to the enhanced dissipation effect.
-To improve it to $\nu^{1/2}$, the key point is the inviscid damping effect and the good structure of the Rayleigh part (or inviscid part)
-$(1-y^{2})+2\Delta^{-1},$
-or in the Fourier version, namely,
-$(1-y^{2})+2(\partial_{y}^{2}-k^{2})^{-1}.$
-The above Rayleigh structure gives good positive estimates for velocity and vorticity(with weight), which along with the diffusion yield better space-time $L^{2}L^{2}$ estimate to vorticity, see Section 2 for instance.
-This implies the good stability mechanism of Poiseuille flow in some sense.
-Moreover, with the good positive structure, one may notice that our proof is different from that of [17] and [14, 18].
-The proof for Theorem 1.2 relies on a direct energy argument with some positive structure for the toy model around Poiseuille flow, see Section 2 for formal analysis.
+Even for the Navier-slip boundary value problem, the transition threshold $\gamma\leq 2/3$ should not be optimal.
+It is very interesting and challenging to study the optimal transition threshold of the Poiseuille flow.
 Let us give some notations to end this section.
@@
 In this section, let us give the sketch and some key ideas in our analysis.
-2.1. The stability for Poiseuille flow with non-slip boundary condition
-To obtain the transition threshold of plane Poiseuille flow with non-slip boundary condition, the key point is to establish the space-time estimates for the following linearized system
-$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\partial_{y}\varphi(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$
-where $f_{i}\in L^{2}$ and $k\not=0$.
-The basic space-time estimates are the $L^{2}L^{2}$ estimates for vorticity and velocity field, which are due to enhanced dissipation and inviscid damping induced by plane Poiseuille flow.
-See Section 3 and [9] for details.
-Compared with the Navier-slip boundary condition in [17], the new effect may result in instability is the boundary layer effect, which is resulted from the non-slip boundary condition.
-Precisely, due to the singularity occurring near the boundaries, there is loss of $\nu$ in the space-time estimates if following the arguments of [17], which will lead to a threshold worse than $\nu^{3/4}$.
-To overcome the difficulty, the main idea is to introduce the suitable weight to control this effect near boundaries $\{y=\pm 1\}$.
-Since near $\{y=\pm 1\}$, it holds that $1-y^{2}\sim 1-|y|$, the suitable weight in our argument is chosen as $1-y^{2}$.
-Precisely, we deduce the weighed estimate as follows:
-$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
-$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2},$
-see Section 3 for details.
-There are some advantages about this weight: it corrects the behavior of the vorticity near boundaries and it can be viewed as the good weight in some Hardy-type inequalities.
-If we consider the estimates without the weight, we only obtain that
-$\nu^{\frac{1}{2}}\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu^{\frac{3}{2}}\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$
-$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2},$
-There is loss of $\nu$ in the above estimates, and the weight $1-y^{2}$ avoid this loss of $\nu$.
-With the above arguments, the space-time estimates for nonlinear stability read as
-$\mathcal{F}_{k}\lesssim\|\omega_{0}\|_{H^{2+}}+\mathrm{nonlinear\ terms},$
-where the energy functional for $k\not=0$ is defined as
-$\mathcal{F}_{k}:=\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}).$
-The term $\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}$ is due to boundary layer effect, $(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}$ is due to enhanced dissipation, and $|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}}$ is due to inviscid damping.
-In other words, the result is the consequence of the balance between the enhanced dissipation, inviscid damping and boundary layer effect.
-Then the nonlinear stability follows from continuous argument, see Subsection 5.1 for details.
-2.2. The improved transition threshold with Navier-slip boundary condition
 In our analysis, compared with [18], the most important point is the improved space-time estimates.
 Indeed, in the view of Fourier transform, the toy model for our problem (linear and nonlinear sense) reads
-$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ u=(\partial_{y},-ik)\varphi,\\ (\varphi,\omega)|_{y=\pm 1}=(0,0),\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (2.1)
-where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$.
+$\left\{\begin{array}[]{lll}\partial_{t}\omega+\mathcal{L}_{k}\omega=-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ u=(\partial_{y},-ik)\varphi,\\ (\varphi,\omega)|_{y=\pm 1}=(0,0),\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (2.1)
+where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$, the linearized operator is defined as
+$\mathcal{L}_{k}:=-\nu(\partial_{y}^{2}-k^{2})+ik(1-y^{2})+2ik(\partial_{y}^{2}-|k|^{2})^{-1}.$
 Let us recall that the space-time estimates in [18] (in the Fourier version) read as
@@
 In a word, the stabilization mechanism used in [18] is only dissipation, including the enhanced dissipation and heat diffusion.
-To establish the improved space-time estimates, we consider the energy form as follows
-$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}$ $\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}$ (2.2)
-$+ik\langle(1-y^{2})\omega+2\varphi,\omega\rangle=\langle-ikf_{1}-\partial_{y}f_{2},\omega\rangle:=\mathrm{RHS}.$
-The first key observation is the inviscid Rayleigh part (imaginary part) of (2.2) will give the space-time estimate for velocity field, which corresponds to the inviscid damping effect (integral form) of velocity field.
-More precisely, the Rayleigh part yields the following positive estimate (see Lemma 4.2)
-$\|(\varphi^{\prime},|k|\varphi)\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\lesssim\langle(1-y^{2})w+2\varphi,w\rangle,$
-which implies the following improved estimate
-$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\left(\|\varphi^{\prime}\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\right)$ (2.3)
-$\lesssim$ $\mathrm{Re}\langle(1-y^{2})\omega+2\varphi,\omega\rangle\lesssim|k|^{-1}|\mathrm{Im}\ \mathrm{RHS}|,$
-which implies the estimate for velocity field. See the proof of Proposition 4.2 for details.
-It should be pointed out that the estimate (2.3) is resulted from the Rayleigh part and it is valid uniform in $\nu$, and this implies the intrinsic structure and stabilization mechanism of the Poiseuille flow.
-The second key observation is the improvement of the $L^{2}L^{2}$-estimate for $\omega$. Precisely, with (2.3) at hands, by applying Lemma 4.1, we have
-$(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}\lesssim(\nu|k|^{2})^{1/3}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{4/3}\|\partial_{y}\omega\|_{L^{2}}^{2/3}\lesssim\nu\|\partial_{y}\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}$
-for $\omega|_{y=\pm 1}=0$. Then we establish
-$\frac{\mathrm{d}}{\mathrm{d}t}$ $\|\omega\|_{L^{2}}^{2}+c(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}+\cdots\lesssim|\mathrm{RHS}|,$ (2.4)
-which suggests that the $L^{2}L^{2}$-estimate for $\omega$ in [17] can be improved as
-$\|e^{c_{0}(\nu|k|^{2})^{1/3}t}\omega\|_{L^{\infty}L^{2}}^{2}+(\nu|k|^{2})^{1/3}\|e^{c_{0}(\nu|k|^{2})^{1/3}t}$ $\omega\|_{L^{2}L^{2}}^{2}+\cdots$ (2.5)
-$\lesssim\|\omega_{0}\|_{L^{2}}^{2}+\mathrm{source\ terms}.$
-Therefore, the improved space-time estimates for toy model (2.1) follow. See Proposition 4.2 for details.
-The nonlinear stability and transition threshold are also based on the improved space-time estimates in Proposition 4.2.
-Indeed, in [17], we only used the enhanced dissipation effect to obtain that transition threshold $\nu^{\gamma}$ with $\gamma\leq 3/4$.
-In this paper, additional to the enhanced dissipation, the inviscid damping for velocity field is established.
-With the effects and Proposition 4.2, the space-time estimates of the $x$-nonzero modes (i.e., in the Fourier version with $k\not=0$) for nonlinear problem read as
-$\mathcal{E}_{k}\lesssim\|\omega_{0,k}\|_{L^{2}}+\mathrm{nonlinear\ terms},$
-where the energy functional $\mathcal{E}_{k}$ with $k\not=0$ is defined as
-$\mathcal{E}_{k}=\|\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}}.$
-Compared with [17], the term $(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}$ is due to the improved space-time estimates, see the formal analysis (2.4)-(2.5).
-If following the arguments of [17], we can only obtain the bound of $(\nu|k|^{2})^{1/4}\|\omega_{k}\|_{L^{2}L^{2}}$, which is worse than the bound obtained in this paper.
-In addition, the term $|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping.
-It should be noted that it is independent of $\nu$ and it would avoid the loss of $\nu$ in nonlinear interaction.
-See Subsection 5.2 for details.
-With the above key points, one can improve the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$.
-3. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow: non-slip boundary condition
+To improve the result of [18] to $\gamma\leq 2/3$, the inviscid damping effect is added in the space-time estimates in this paper.
+More precisely, the following space-time estimates are used in this paper
+$\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$
+$+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c\nu^{\frac{1}{2}}t}(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
+$+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|e^{c\nu^{\frac{1}{2}}t}f_{4}\|_{L^{2}L^{2}}^{2}.$
+Compared with the space-time estimates in [18], in the above estimates, the part $|k|\|e^{c\nu^{\frac{1}{2}}t}u\|_{L^{2}L^{2}}^{2}$ is due to the inviscid damping effect of the Poiseuille flow.
+To achieve the desired space-time estimates, following the idea of [9], the key point is to establish the resolvent estimates for the linearized operator around the Poiseuille flow, see Section 3 for details.
+With the resolvent estimates at hands, one can achieve the space-time estimates for the linearized N-S system.
+For the nonlinear problem, the energy estimates for non-zero modes read as
+$\mathcal{E}_{k}\lesssim\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}},$
+where $\mathcal{E}_{k}$ is defined as
+$\mathcal{E}_{k}:=$ $\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}+\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$
+$+(\nu|k|)^{1/4}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}}.$
+In the above functional, the terms $\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}$ and $\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$ are corresponding to the basic energy and heat diffusion, the term $(\nu|k|)^{1/4}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$ is due to the enhanced dissipation, and $|k|^{1/2}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping effect of Poiseuille flow.
+With the nonlinear estimates, the nonlinear stability follows from a continuous argument.
+3. Resolvent estimates for Orr-Sommerfeld equation
+In this section, we study the resolvent estimates for the Orr-Sommerfeld equation around the plane Poiseuille flow as follows
+$\left\{\begin{array}[]{lll}-\nu(\partial_{y}^{2}-|k|^{2})w+ik[(1-y^{2}-\lambda)w+2\varphi]=F,\\ (\partial_{y}^{2}-|k|^{2})\varphi=w,\ \varphi(\pm 1)=w(\pm 1)=0,\end{array}\right.$ (3.1)
+where $k\neq 0$ and $F\in H^{1}_{0}$.
+The resolvent estimates for (3.1) are stated as follows.
+Proposition 3.1.
+Let $w$ be the solution to (3.1), then it holds that
+$\nu^{\frac{3}{8}}$ $|k|^{\frac{5}{8}}(\|w\|_{L^{1}}+|k|^{\frac{1}{2}}\|u\|_{L^{2}})$ (3.2)
+$+(\nu|k|)^{\frac{1}{2}}\|w\|_{L^{2}}+\nu^{\frac{3}{4}}|k|^{\frac{1}{4}}\|(\partial_{y},|k|)w\|_{L^{2}}\lesssim\|F\|_{L^{2}},$
+$\nu^{\frac{3}{4}}$ $|k|^{\frac{1}{4}}\|w\|_{L^{2}}+\nu\|(\partial_{y},|k|)w\|_{L^{2}}+(\nu|k|)^{\frac{1}{2}}\|u\|_{L^{2}}\lesssim\|F\|_{H^{-1}_{k}},$
+||| $\nu/k|^{\frac{1}{2}}\|(\partial_{y},|k|)w\|_{L^{2}}+|\nu/k|^{\frac{1}{4}}\|w\|_{L^{2}}$
+$+|\nu/k|^{\frac{1}{8}}\|u\|_{L^{\infty}}+\|u\|_{L^{2}}\lesssim|k|^{-1}\|(\partial_{y},|k|)F\|_{L^{2}}.$
+Moreover, there holds that
+$\nu^{\frac{1}{6}}$ $|k|^{\frac{5}{6}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{1}{3}}\|u\|_{L^{2}}$ (3.3)
+$+\nu^{\frac{2}{3}}|k|^{\frac{1}{3}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{1}{3}}\|(\partial_{y},|k|)w\|_{L^{2}}$
+$+\nu^{\frac{1}{3}}|k|^{\frac{2}{3}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{2}{3}}\|w\|_{L^{2}}\lesssim\|F\|_{L^{2}},$
+$\nu^{\frac{2}{3}}$ $|k|^{\frac{1}{3}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{1}{3}}\|w\|_{L^{2}}\lesssim\|F\|_{H^{-1}_{k}}.$
+Here, $\|F\|_{H^{-1}_{k}}:=\inf\limits_{\{g\in H^{1}_{0}:\|g\|_{H^{1}_{k}}=1\}}\langle f,g\rangle$ with $\|g\|_{H^{1}_{k}}=\|(\partial_{y},|k|)g\|_{L^{2}}.$
+Proof.
+See [9]. ∎
+4. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow
 The linearized Navier-Stokes equations around Poiseuille flow read as
-$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\partial_{y}\varphi(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (3.1)
-where $f_{i}\in L^{2}$ and $k\not=0$.
+$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\omega(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (4.1)
+where $f_{3}|_{y=\pm 1}=0$ and $k\not=0$.
 The space-time estimates for (4.1) are stated as follows.
 Proposition 4.1.
-Suppose the $\omega_{0}\in H^{1}_{0}$ with $\langle\omega_{0},e^{\pm|k|y}\rangle=0$, then the solution $\omega$ to (3.1) enjoys that the following estimates:
-(1) If $\nu|k|^{2}\leq 2$, it holds that
-$(\nu|k|)^{\frac{1}{2}}$ $\|\omega\|_{L^{2}L^{2}}^{2}+|k|(\|u\|_{L^{\infty}L^{2}}^{2}+\|u\|_{L^{2}L^{2}}^{2})+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}L^{2}}^{2}$ (3.2)
-$+\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
-$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
-(2) If $\nu|k|^{2}\geq 2$, it holds that
-$\|\omega\|_{L^{\infty}L^{2}}^{2}$ $+|k|^{2}(\|u\|_{L^{\infty}L^{2}}^{2}+\|u\|_{L^{2}L^{2}}^{2})+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$ (3.3)
-$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
+Suppose the $\omega_{0}\in H^{1}_{0}\cap H^{2}$, then the solution $\omega$ to (4.1) enjoys that the following estimates:
+$\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega\|_{L^{2}L^{2}}^{2}+|k|\|u\|_{L^{2}L^{2}}^{2}$ (4.2)
+$\lesssim$ $\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}$
+$+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
+Moreover, there exists $0<c\ll 1$ such that
+$\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$ (4.3)
+$+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c\nu^{\frac{1}{2}}t}(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
+$+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|e^{c\nu^{\frac{1}{2}}t}f_{4}\|_{L^{2}L^{2}}^{2}.$
+To establish Proposition 4.1, let us introduce
+$\omega=\omega_{I}+\omega_{H}$
+with
+$\left\{\begin{array}[]{lll}\partial_{t}\omega_{I}-\nu(\partial_{y}^{2}-k^{2})\omega_{I}+ik(1-y^{2})\omega_{I}+2ik\varphi_{I}=-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi_{I}=\omega_{I},\ \ u_{I}=(\partial_{y},-ik)\varphi_{I},\\ \varphi_{I}(t,k,\pm 1)=\omega_{I}(t,k,\pm 1)=0,\\ \omega_{I}|_{t=0}=0,\end{array}\right.$ (4.4)
+and
+$\left\{\begin{array}[]{lll}\partial_{t}\omega_{H}-\nu(\partial_{y}^{2}-k^{2})\omega_{H}+ik(1-y^{2})\omega_{H}+2ik\varphi_{H}=0,\\ (\partial_{y}^{2}-k^{2})\varphi_{H}=\omega_{H},\ \ u_{H}=(\partial_{y},-ik)\varphi_{H},\\ \varphi_{H}(t,k,\pm 1)=\omega_{H}(t,k,\pm 1)=0,\\ \omega_{H}|_{t=0}=\omega_{0}.\end{array}\right.$ (4.5)
+The space-time estimates for $\omega_{I}$ are stated as follows.
+Lemma 4.1.
+Let $\omega_{I}$ be the solution to (4.4), then it holds that
+$\|\omega_{I}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}+|k|\|u_{I}\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
+In addition, there exists $0<c\ll 1$ such that
+$\|e^{c\nu^{\frac{1}{2}}t}\omega_{I}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{I}\|_{L^{2}L^{2}}^{2}$ (4.6)
+$\qquad+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u_{I}\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|e^{c\nu^{\frac{1}{2}}t}(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}$
+$+|k|^{-1}\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|e^{c\nu^{\frac{1}{2}}t}f_{4}\|_{L^{2}L^{2}}^{2}.$
 Proof.
-Indeed, it remains to give the space-time estimates of $\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}$ and $|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$ for $\nu|k|^{2}\leq 2$, other estimates can be found in [9].
-By testing (3.1) by $(1-y^{2})\omega$, we have
-$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+\nu\int_{-1}^{1}(-2y)\partial_{y}\omega\overline{\omega}\mathrm{d}y$
-$+ik\|(1-y^{2})\omega\|_{L^{2}}^{2}=\langle-ikf_{1},\omega\rangle+\langle f_{2},\partial_{y}((1-y^{2})\omega)\rangle+2ik\langle\varphi,(1-y^{2})\omega\rangle,$
+Note that the proof is based on the resolvent estimates for $\mathcal{L}_{k}-\epsilon(\nu|k|)^{\frac{1}{2}}$, we only give the proof for the estimates without $e^{c\nu^{\frac{1}{2}}t}$, and the case with the weight can be obtained via perturbation argument.
+By Proposition 3.1, we have
+$(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}$
+$+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2},$
+where the following energy estimates have been used
+$\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+(\nu|k|^{2})^{-1}\|f_{4}\|_{L^{2}L^{2}}^{2}$
+and
+$\nu|k|^{2}\|\omega_{I}\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+\|f_{4}\|_{L^{2}L^{2}}\|\omega\|_{L^{2}L^{2}}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+(\nu|k|)^{-\frac{1}{2}}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
+It remains to bound $\|u_{I}\|_{L^{2}L^{2}}$ and $\|\omega_{I}\|_{L^{\infty}L^{2}}$. Let us introduce
+$\omega_{I}=\omega_{I}^{(1)}+\omega_{I}^{(2)}$
+with
+$\left\{\begin{array}[]{lll}\partial_{t}\omega_{I}^{(1)}-\nu(\partial_{y}^{2}-k^{2})\omega_{I}^{(1)}+ik(1-y^{2})\omega_{I}^{(1)}+2ik\varphi_{I}^{(1)}=-ikf_{1}-\partial_{y}f_{2}-f_{3},\\ \partial_{t}\omega_{I}^{(2)}-\nu(\partial_{y}^{2}-k^{2})\omega_{I}^{(2)}+ik(1-y^{2})\omega_{I}^{(2)}+2ik\varphi_{I}^{(2)}=f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi_{I}^{(i)}=\omega_{I}^{(i)},\ \ u_{I}^{(i)}=(\partial_{y},-ik)\varphi_{I}^{(i)},\\ \varphi_{I}^{(i)}(t,k,\pm 1)=\omega_{I}^{(i)}(t,k,\pm 1)=0,\ \omega_{I}^{(i)}|_{t=0}=0,\end{array}\right.$ (4.7)
+By Proposition 3.1, we have
+$|k|\|u_{I}^{(1)}\|_{L^{2}L^{2}}^{2}\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}$
+and
+$|k|\|u_{I}^{(2)}\|_{L^{2}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim(\nu|k|)^{-\frac{1}{2}}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
+Meanwhile, one has
+$|k|\|u_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\leq|k|^{-1}\|\omega_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim(\nu|k|^{2})^{-1}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
+Thus, it holds that
+$|k|\|u_{I}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
+Putting the above estimates together, the estimate for $\|u_{I}\|_{L^{2}L^{2}}$ follows.
+Finally, we give the estimate for $\|\omega_{I}\|_{L^{\infty}L^{2}}$. Indeed, we have
+$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{I}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}=\mathrm{Re}\langle-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\omega_{I}\rangle$
+$\leq$ $\|(f_{1},f_{2})\|_{L^{2}}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}+\|(\partial_{y},|k|)f_{3}\|_{L^{2}}\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}+\|f_{4}\|_{L^{2}}\|\omega_{I}\|_{L^{2}}$
+$\leq$ $\frac{1}{2}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+|k|\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}^{2}$
+$+(\nu|k|^{2})^{-1}\|f_{4}\|_{L^{2}}^{2}$
+or
+$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{I}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}=\mathrm{Re}\langle-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\omega_{I}\rangle$
+$\leq$ $\|(f_{1},f_{2})\|_{L^{2}}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}+\|(\partial_{y},|k|)f_{3}\|_{L^{2}}\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}+\|f_{4}\|_{L^{2}}\|\omega_{I}\|_{L^{2}}$
+$\leq$ $\frac{1}{2}\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}+|k|\|(\partial_{y},|k|)\varphi_{I}\|_{L^{2}}^{2}$
+$+\frac{1}{2}\left((\nu|k|)^{-1/2}\|f_{4}\|_{L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}}^{2}\right),$
 which gives that
-$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}$
-$\leq$ $2\nu\|\omega\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}+2|k|\|\varphi\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}$
-$+|k|\|f_{1}\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}((1-y^{2})\omega)\|_{L^{2}}$
-$\leq$ $\nu^{\frac{1}{2}}\|\omega\|_{L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}+4|k|\|\varphi\|_{L^{2}}^{2}$
-$+\frac{1}{2}\nu|k|^{2}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\frac{1}{2}(\nu|k|^{2})^{-1}\|(1-y^{2})^{\frac{1}{2}}f_{1}\|_{L^{2}}^{2}$
-$+\frac{1}{2}\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{2}\nu\|(1-y^{2})^{\frac{1}{2}}\partial_{y}\omega\|_{L^{2}}^{2}$
-or
-$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}$
-$\leq$ $2\nu\|\omega\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}+2|k|\|\varphi\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}$
-$+|k|\|f_{1}\|_{L^{2}}\|(1-y^{2})\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}((1-y^{2})\omega)\|_{L^{2}}$
-$\leq$ $\nu^{\frac{1}{2}}\|\omega\|_{L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})\omega\|_{L^{2}}^{2}+4|k|\|\varphi\|_{L^{2}}^{2}$
-$+\frac{1}{2}|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\frac{1}{2}|k|\|(1-y^{2})^{\frac{1}{2}}f_{1}\|_{L^{2}}^{2}$
-$+\frac{1}{2}\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{2}\nu\|(1-y^{2})^{\frac{1}{2}}\partial_{y}\omega\|_{L^{2}}^{2}.$
-Therefore, we obtain that
-$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
-$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\nu^{\frac{1}{2}}\|\omega\|_{L^{2}L^{2}}^{2}+\nu^{\frac{3}{2}}\|\partial_{y}\omega\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|u\|_{L^{2}L^{2}}^{2}$
-$+\min\{(\nu|k|^{2})^{-1},|k|\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}$
-$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}.$
+$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{I}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}}^{2}$
+$+\min\{(\nu|k|)^{-1/2},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}}^{2}+|k|\|u_{I}\|_{L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}}^{2}.$
+Therefore, we have
+$\|\omega_{I}\|_{L^{\infty}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{I}\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
+$+\min\{(\nu|k|)^{-1/2},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}+|k|\|u_{I}\|_{L^{2}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{I}\|_{L^{2}L^{2}}^{2}$
+$\lesssim$ $\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
+$+\min\{(\nu|k|)^{-1/2},(\nu|k|^{2})^{-1}\}\|f_{4}\|_{L^{2}L^{2}}^{2}.$
 The proof is completed. ∎
-4. Space-time estimates for linearized Navier-Stokes equations around the Poiseuille flow: Navier-slip boundary condition
-In this section, we will establish the space-time estimates for linearized Navier-Stokes equations as follows
-$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\omega(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (4.1)
-where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$. It is noted that $\omega_{0}(k,\pm 1)=0$ ensures $\omega(t,k,\pm 1)=0$ for any $t\geq 0$.
-Before giving the main result of this section, let us recall two lemmas from [9].
+Next we give the estimates for $\omega_{H}$.
 Lemma 4.2.
-For $f\in H^{1}_{0}(-1,1)$, it holds that
-$\|f\|_{L^{2}}^{2}\leq\|(1-y^{2})^{\frac{1}{2}}f\|_{L^{2}}^{\frac{4}{3}}\|(1-y^{2})^{-1}f\|_{L^{2}}^{\frac{2}{3}}\leq C\|(1-y^{2})^{\frac{1}{2}}f\|_{L^{2}}^{\frac{4}{3}}\|\partial_{y}f\|_{L^{2}}^{\frac{2}{3}}.$
-Lemma 4.2.
-Suppose that $(\partial_{y}^{2}-|k|^{2})\varphi=w,\varphi(\pm 1)=0,|k|\neq 0$. Then it holds that
-$2\left(\bigg{\|}\varphi^{\prime}+\dfrac{2y\varphi}{1-y^{2}}\bigg{\|}_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\right)\leq\langle(1-y^{2})w+2\varphi,w\rangle,$
-$\|(\varphi^{\prime},|k|\varphi)\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\leq\big{(}8/|k|^{2}+1\big{)}\langle(1-y^{2})w+2\varphi,w\rangle.$
-The space-time estimates for (4.1) are stated as follows.
-Proposition 4.1.
-Let $\omega$ be the solution to (4.1) with $\omega_{0}\in H^{1}_{0}(-1,1)$, and $f_{i}\in L^{2}(i=1,2)$, then there exists $0<c_{0}\ll 1$ such that
-$\|\omega\|_{L^{\infty}L^{2}}^{2}+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$ (4.2)
-$\qquad+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}L^{2}}^{2}$
-$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}$
+It holds that
+[…]
+$+(\nu|k|)^{\frac{1}{2}}\|\omega_{H}\|_{L^{2}L^{2}}^{2}+|k|\|u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}.$
+In addition, there exists $0<c\ll 1$ such that
+$\|e^{c\nu^{\frac{1}{2}}t}$ $\omega_{H}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{H}\|_{L^{2}L^{2}}^{2}$
+$+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega_{H}\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}.$
+Proof.
+The proof is based on the resolvent estimates for $\mathcal{L}_{k}-\epsilon(\nu|k|)^{\frac{1}{2}}$, we only give the proof for the estimates without $e^{c\nu^{\frac{1}{2}}t}$, and the case with the weight can be obtained via perturbation argument with standard energy argument.
+By the Gearhart-Prüss type lemma in [43] and the resolvent estimates in Proposition 3.1, we have
+$\|\omega_{H}(t)\|_{L^{2}}\leq Ce^{-c_{1}(\nu|k|)^{\frac{1}{2}}t}\|\omega_{0}\|_{L^{2}},$
+which gives that
+$\|\omega_{H}\|_{L^{\infty}L^{2}}^{2}+(\nu|k|)^{\frac{1}{2}}\|\omega_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\omega_{0}\|_{L^{2}}^{2}.$
+The energy argument gives that
+$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega_{H}\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega_{H}\|_{L^{2}}^{2}=0,$
+which yields that
+$\nu\|(\partial_{y},|k|)\omega_{H}\|_{L^{2}L^{2}}^{2}\leq\|\omega_{0}\|_{L^{2}}^{2}.$
+Next we give the estimate for $u_{H}$.
+If $\nu|k|^{2}\leq 2$, let $\omega_{H}^{(0)}$ be the solution to
+$\partial_{t}\omega_{H}^{(0)}+ik(1-y^{2})\omega_{H}^{(0)}+2ik\varphi_{H}^{(0)}=0,$
+$(\partial_{y}^{2}-|k|^{2})\varphi_{H}^{(0)}=\omega_{H}^{(0)},\varphi_{H}^{(0)}|_{y=\pm 1}=0,\ \omega_{H}^{(0)}|_{t=0}=\omega_{0},$
+due to $\omega_{0}\in H^{1}_{0}$, it holds that $\omega_{H}^{(0)}|_{y=\pm 1}=0$ for any $t\geq 0$.
+Following the arguments of [9], let us introduce
+$\omega_{H}^{(1)}=g(t,y)\omega^{(0)}_{H},\ g(t,y)=e^{-\nu|k|^{2}t-4\nu|k|^{2}y^{2}t^{3}/3-(\nu|k|)^{1/2}t},$
+$(\partial_{y}^{2}-|k|^{2})\varphi_{H}^{(i)}=\omega_{H}^{(i)},\quad\varphi_{H}^{(i)}(\pm 1)=0,\quad i=0,1.$
+It is easy to find that
+$\partial_{t}\omega_{H}^{(1)}+\nu|k|^{2}\omega_{H}^{(1)}+(4\nu|k|^{2}y^{2}t^{2}+(\nu|k|)^{1/2})\omega_{H}^{(1)}$
+$\qquad\qquad+ik(1-y^{2})\omega_{H}^{(1)}=-2ikg(t,y)\varphi_{H}^{(0)},$
+$\omega_{H}^{(0)}|_{t=0}=\omega_{H}^{(1)}|_{t=0}=\omega_{0},\ \omega_{H}^{(1)}|_{y=\pm 1}=0.$
+Thus, we have
+$\partial_{t}\omega_{H}^{(1)}-\nu(\partial_{y}^{2}-|k|^{2})\omega_{H}^{(1)}+ik(1-y^{2})\omega_{H}^{(1)}+2ik\varphi_{H}^{(1)}$
+$=-\nu\partial_{y}^{2}\omega_{H}^{(1)}-4\nu|k|^{2}y^{2}t^{2}\omega_{H}^{(1)}-(\nu|k|)^{1/2}\omega_{H}^{(1)}-2ik(g\varphi_{H}^{(0)}-\varphi_{H}^{(1)}).$
+Now we decompose $\omega_{H}$ as follows
+$\omega_{H}=\omega_{H}^{(1)}+\omega_{H}^{(2)},$
+where $\omega_{H}^{(2)}$ satisfies
+$\left\{\begin{array}[]{lll}(\partial_{t}+\mathcal{L}_{k})\omega_{H}^{(2)}=\nu\partial_{y}^{2}\omega_{H}^{(1)}+4\nu|k|^{2}y^{2}t^{2}\omega_{H}^{(1)}+|\nu k|^{1/2}\omega_{H}^{(1)}\\ \qquad\qquad\qquad\quad+2ik(g(t,y)\varphi_{H}^{(0)}-\varphi_{H}^{(1)}),\\ (\partial_{y}^{2}-|k|^{2})\varphi_{H}^{(2)}=\omega_{H}^{(2)},\ \omega_{H}^{(2)}|_{t=0}=0,\ \omega_{H}^{(2)}|_{y=\pm 1}=0,\end{array}\right.$
+Following the arguments as in Lemma 4.2 of [9], we have
+$|k|\|(\partial_{y},|k|)\varphi_{H}^{(1)}\|_{L^{2}L^{2}}^{2}\lesssim\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}.$ (4.8)
+Next, we estimate $\|(\partial_{y},|k|)\varphi_{H}^{(2)}\|_{L^{2}L^{2}}$. Let us recall from Lemma 4.2 in [9] (here we set $k_{1}=k$) that
+$\partial_{y}^{2}\omega^{(1)}_{H}$ $=\partial_{y}\omega_{1}^{(1)}+2ikt\omega_{H}^{(1)}+2ikty(\omega_{1}^{(1)}+2ikty\omega^{(1)}_{H}),$
+$\partial_{y}^{2}\omega_{H}^{(1)}$ $+4|k|^{2}y^{2}t^{2}\omega^{(1)}_{H}=\partial_{y}\omega_{1}^{(1)}+2ikt\omega^{(1)}_{H}+2ikty\omega_{1}^{(1)},$
+$\omega_{1}=X{\omega}^{(0)}_{H},\ \omega_{1}^{(1)}=X\omega^{(1)}_{H},\ X=\partial_{y}-2ikyt.$
+$\phi_{1}^{(1)}=g(t,y)\varphi_{H}^{(0)}-\varphi_{H}^{(1)},$
+and we have
+$\partial_{t}\omega^{(2)}_{H}-\nu(\partial_{y}^{2}-|k|^{2})\omega_{H}^{(2)}+ik\big{(}(1-y^{2})\omega_{H}^{(2)}+2\varphi^{(2)}_{H}\big{)}$
+$=\nu(\partial_{y}\omega_{1}^{(1)}+2ikt\omega^{(1)}_{H}+2ikty\omega_{1}^{(1)})+|\nu k|^{1/2}\omega_{H}^{(1)}+2ik\phi_{1}^{(1)}.$
+Applying Lemma 4.1 with $f_{2}=-\nu\omega_{1}^{(1)}$, $f_{4}=-2\nu ikty\omega_{1}^{(1)}-2\nu ikt\omega_{H}^{(1)}-(\nu|k|)^{\frac{1}{2}}\omega_{H}^{(1)}$ and $f_{3}=-2ik\phi_{1}^{(1)}$ that
+$\|(\partial_{y},|k|)\varphi_{H}^{(2)}\|_{L^{2}L^{2}}^{2}\lesssim$ $\nu^{-1}|k|^{-1}\|\nu\omega_{1}^{(1)}\|_{L^{2}L^{2}}^{2}+|k|^{-2}\|k(\partial_{y},|k|)\phi_{1}^{(1)}\|_{L^{2}L^{2}}^{2}$
+$+\nu^{-\frac{1}{2}}|k|^{-\frac{3}{2}}\bigg{(}\|\nu kty\omega_{1}^{(1)}\|_{L^{2}L^{2}}^{2}+\|(|\nu kt|+(\nu|k|)^{\frac{1}{2}})\omega_{H}^{(1)}\|_{L^{2}L^{2}}^{2}\bigg{)}$
+$\lesssim$ $|k|^{-1}\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}.$
+Therefore, we conclude that
+$|k|\|u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2},$
+which completes the proof for case of $\nu|k|^{2}\leq 2$.
+If $\nu|k|^{2}\geq 2$, then it holds that
+$|k|^{2}\|u_{H}\|_{L^{2}L^{2}}^{2}\lesssim\nu|k|^{2}\|\omega_{H}\|_{L^{2}L^{2}}^{2}\lesssim\|\omega_{0}\|_{L^{2}}^{2},$
+which completes that proof. ∎
+5. Nonlinear stability
+For the nonlinear problem with Navier-slip boundary condition, let us recall that
+$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}+\mathcal{L}_{k}\omega_{k}=-ik(f_{1})_{k}-\partial_{y}(f_{2})_{k},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\\ \varphi_{k}(\pm 1)=\omega_{k}(\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.1)
+where
+$\mathcal{L}_{k}:=-\nu(\partial_{y}^{2}-k^{2})+ik(1-y^{2})+2ik(\partial_{y}^{2}-|k|^{2})^{-1}$
 and
-$\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{\infty}L^{2}}^{2}$ (4.3)
-$\qquad+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}L^{2}}^{2}$
-$\qquad+|k|\left(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
-$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}L^{2}}^{2}.$
-Proof.
-The standard energy arguments give that
-$\frac{1}{2}$ $\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+ik\langle(1-y^{2})\omega+2\varphi,\omega\rangle$
-$=$ $\langle-ikf_{1},\omega\rangle+\langle f_{2},\partial_{y}\omega\rangle=:\mathrm{RHS}.$
-By Lemma 4.2, we have
-$|k|$ $\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}\lesssim|\mathrm{Im}\ \mathrm{RHS}|,$
-which gives that
-$|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}\lesssim|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}+|\mathrm{Im}\ \mathrm{RHS}|\lesssim|\mathrm{Im}\ \mathrm{RHS}|.$
-Therefore, we have
-$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\frac{\nu}{2}\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}}^{2}$
-$\qquad\qquad+\frac{|k|}{2}(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
-$\leq$ $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}$
-$\qquad+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}\right)\leq C|\mathrm{RHS}|,$
-where we have used Lemma 4.1 for $\omega(\pm 1)=0$.
-The RHS is bounded as
-$|\mathrm{RHS}|\leq$ $|k|\|f_{1}\|_{L^{2}}\|\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}$
-$\leq$ $\frac{1}{4}\left(\nu|k|^{2}\|\omega\|_{L^{2}}^{2}+\nu\|\partial_{y}\omega\|_{L^{2}}^{2}\right)+C\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}}^{2}+\frac{1}{4}|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2},$
-or
-$|\mathrm{RHS}|\leq$ $|k|\|f_{1}\|_{L^{2}}\|\omega\|_{L^{2}}+\|f_{2}\|_{L^{2}}\|\partial_{y}\omega\|_{L^{2}}$
-$\leq$ $\epsilon_{1}(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}+C(\nu|k|^{2})^{-1/3}|k|^{2}\|f_{1}\|_{L^{2}}^{2}$
-$+\frac{1}{4}\nu\|\partial_{y}\omega\|_{L^{2}}^{2}+C\nu^{-1}\|f_{2}\|_{L^{2}}^{2}+\frac{1}{4}|k|\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2}.$
-Taking $\epsilon_{1}\ll\min\{c,1\}$, we have
-$\frac{\mathrm{d}}{\mathrm{d}t}\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}}^{2}$
-$\qquad\qquad\qquad+|k|(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
-$\lesssim$ $\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}}^{2}.$
-Therefore, we get that
-$\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|\omega\|_{L^{2}L^{2}}^{2}$
-$\qquad\qquad\qquad+|k|\left(\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
-$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|f_{2}\|_{L^{2}L^{2}}^{2}.$
-Moreover, we have
-$\frac{\mathrm{d}}{\mathrm{d}t}$ $\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}$
-$+|k|(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}}^{2})$
-$\lesssim$ $\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}}^{2}$
-$+c_{0}(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}}^{2}.$
-Taking $c_{0}\ll 1$, we have
-$\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+c(\nu k^{2})^{\frac{1}{3}}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}\omega\|_{L^{2}L^{2}}^{2}$
-$+|k|\left(\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}L^{2}}^{2}+\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}(\partial_{y},|k|)\varphi\|_{L^{2}L^{2}}^{2}\right)$
-$\lesssim$ $\|\omega_{0}\|_{L^{2}}^{2}+\min\{(\nu^{-1}k^{4})^{\frac{1}{3}},\nu^{-1}\}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{1}\|_{L^{2}L^{2}}^{2}+\nu^{-1}\|e^{c_{0}(\nu k^{2})^{\frac{1}{3}}t}f_{2}\|_{L^{2}L^{2}}^{2}.$
-The proof is completed. ∎
-5. Nonlinear stability
-In this section, we will give the proof of the main results, i.e., Theorem 1.1 and Theorem 1.2.
-5.1. Nonlinear stability with non-slip boundary condition: Proof of Theorem 1.1
-Recall that the Navier-Stokes equations around Poiseuille flow with non-slip boundary condition read as
-$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}-\nu(\partial_{y}^{2}-k^{2})\omega_{k}+ik(1-y^{2})\omega_{k}+2ik\varphi_{k}=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\ \ u_{k}=(\partial_{y},-ik)\varphi_{k},\\ \varphi_{k}(t,k,\pm 1)=\partial_{y}\varphi_{k}(t,k,\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.1)
-where
-$(f_{1})_{k}=-\sum_{l\in\mathbb{Z}}(u_{1})_{l}\omega_{k-l},\ (f_{2})_{k}=-\sum_{l\in\mathbb{Z}}(u_{2})_{l}\omega_{k-l}.$
-Let us recall that
-$\mathcal{F}_{k}:=\left\{\begin{array}[]{lll}\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}\\ \qquad\qquad\qquad+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}),&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$
-Now we are ready to prove Theorem 1.1.
-Proof of Theorem 1.1.
-The global well-posedness of 2-D Navier-Stokes equation around the Poiseuille flow can be obtained by classical theory for Navier-Stokes equation.
-Here we only focus on the stability estimates.
-It should be noted that the case of $\nu|k|^{2}\geq 2$ is trivial by using (3.3) and we only consider the case of $\nu|k|^{2}\leq 2$.
-Let us first consider the zero modes.
-If $k=0$, by the divergence-free condition, we have $\partial_{y}(u_{2})_{0}=0$, which implies by $(u_{2})_{0}|_{y=\pm 1}=0$ that $(u_{2})_{0}=0$.
-Therefore, the equation for zero modes is reduced to
-$\partial_{t}(u_{1})_{0}-\nu\partial_{y}^{2}(u_{1})_{0}=-\partial_{y}(f_{2})_{0}.$
-By the similar arguments as for 2-D Couette flow [10], we have
-$\mathcal{F}_{0}^{2}\lesssim\nu^{-1}\|(f_{2})_{0}\|_{L^{2}L^{2}}^{2}+\|\overline{\omega_{0}}\|_{L^{2}}^{2}.$ (5.2)
-For $\mathcal{F}_{k}$, by Proposition 3.1, one has
-$\mathcal{F}_{k}\lesssim|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}.$
-For $k\not=0$, it holds that
-$\|(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$ (5.3)
-$\leq$ $\sum_{l\in\mathbb{Z}}\left\|(u_{2})_{l}\right\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}$
-$\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\|\partial_{y}(u_{2})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
-$\lesssim$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}(|l|\|(u_{1})_{l}\|_{L^{\infty}L^{2}})^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
-$\lesssim$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}(|l|^{\frac{1}{2}}\mathcal{F}_{l})^{\frac{1}{2}}\nu^{-\frac{1}{4}}|k-l|^{-\frac{1}{4}}\mathcal{F}_{k-l}$
-$\lesssim$ $\nu^{-\frac{1}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}$
-and
-$\|(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.4)
-$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|(u_{1})_{l}\omega_{k-l}\|_{L^{2}L^{2}}.$
-It remains to estimate the terms on the right-hand side of (5.9).
-For $y\in[-1,0]$, it holds that $(1-|y|)=(1+y)\leq(1+y)(1-y)=1-y^{2}$; and for $y\in[0,1]$, $(1-|y|)=(1-y)\leq(1+y)(1-y)=1-y^{2}$, thus $1-y^{2}\geq 1-|y|$.
-For $l\not=0,k$, it holds that $|k|\lesssim|l||k-l|$, then
-$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|(u_{1})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$
-$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\left\|\frac{(u_{1})_{l}}{(1-y^{2})^{\frac{1}{2}}}\right\|_{L^{2}L^{\infty}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
-$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|\partial_{y}(u_{1})_{l}\|_{L^{2}L^{2}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
-$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|\omega_{l}\|_{L^{2}L^{2}}\|(1-y^{2})^{\frac{1}{2}}\omega_{k-l}\|_{L^{\infty}L^{2}}$
-$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}(\nu|l|)^{-\frac{1}{4}}\mathcal{F}_{l}\mathcal{F}_{k-l}$
-and
-$\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}$ $\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
-$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}\|\omega_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{4}}|k|^{-\frac{1}{4}}\mathcal{F}_{0}\mathcal{F}_{k}.$
-Thus, we have
-$\|(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{4}}|k|^{-\frac{1}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}$ (5.5)
-Combining (5.2), (5.3) and (5.5), we have
-$\mathcal{F}_{k}\lesssim$ $|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}$
-$\lesssim$ $|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{4}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l},$
-$\mathcal{F}_{0}\lesssim$ $\|\overline{\omega}_{0}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}\setminus\{0\}}\mathcal{F}_{l}\mathcal{F}_{-l},$
-which gives that
-$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\lesssim\sum_{k\in\mathbb{Z}}|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{4}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{F}_{l}\mathcal{F}_{k-l}.$
-Due to $\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{\frac{3}{4}}$, then it holds that
-$\sum_{k\in\mathbb{Z}}|k|^{-\frac{1}{2}}\|(\partial_{y},|k|)\omega_{0,k}\|_{L^{2}}\leq\epsilon_{0}C\nu^{\frac{3}{4}}.$
-If $\epsilon_{0}$ is suitably small, one can deduce from continuous argument that
-$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\leq\epsilon_{0}C\nu^{\frac{3}{4}}.$
-The proof is completed. ∎
-5.2. Nonlinear stability with Navier-slip boundary condition: Proof of Theorem 1.2
-For the nonlinear problem with Navier-slip boundary condition, let us recall that
-$\left\{\begin{array}[]{lll}\partial_{t}\omega_{k}-\nu(\partial_{y}^{2}-k^{2})\omega_{k}+ik(1-y^{2})\omega_{k}+2ik\varphi_{k}=-ik(f_{1})_{k}-\partial_{y}(f_{2})_{k},\\ (\partial_{y}^{2}-k^{2})\varphi_{k}=\omega_{k},\ \varphi_{k}(\pm 1)=\omega_{k}(\pm 1)=0,\\ \omega_{k}|_{t=0}=\omega_{0,k}(k,y),\end{array}\right.$ (5.6)
-where
 $(f_{1})_{k}=-\sum_{l\in\mathbb{Z}}(u_{1})_{l}\omega_{k-l},\ (f_{2})_{k}=-\sum_{l\in\mathbb{Z}}(u_{2})_{l}\omega_{k-l}.$
@@
 Recall that for $k\not=0,$
-$\mathcal{E}_{k}=(\nu|k|^{2})^{\frac{1}{6}}\|\omega_{k}\|_{L^{2}L^{2}}+\|\omega_{k}\|_{L^{\infty}L^{2}}+|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}},$
+$\mathcal{E}_{k}=$ $(\nu|k|)^{\frac{1}{4}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$
+$+\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}+|k|^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}},$
 and $\mathcal{E}_{0}=\|\overline{\omega}\|_{L^{\infty}L^{2}}$ for $k=0$.
@@
 For $\mathcal{E}_{k}$, by the space-time estimates in Proposition 4.1, one has
-$\mathcal{E}_{k}\lesssim\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{3}}|k|^{\frac{1}{3}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}},$
-where we have used the fact that $\min\{(\nu^{-1}|k|^{4})^{\frac{1}{3}},\nu^{-1}\}\leq(\nu|k|^{-1})^{-\frac{2}{3}}$.
+$\mathcal{E}_{k}\lesssim\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}}.$
+where we have used the fact that $|k|\min\{(\nu|k|)^{-\frac{1}{4}},(\nu|k|^{2})^{-\frac{1}{2}}\}\leq\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}$.
 For $k\not=0$, it holds that
-$\|(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}\leq\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{2}L^{\infty}}\|\omega_{k-l}\|_{L^{\infty}L^{2}}$ (5.8)
-$\leq$ $\sum_{l\in\mathbb{Z}}\|(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|\partial_{y}(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\mathcal{E}_{k-l}$
-$\leq$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{E}_{l})^{\frac{1}{2}}(|l|\|(u_{1})_{l}\|_{L^{2}L^{2}})^{\frac{1}{2}}\mathcal{E}_{k-l}$
+$\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}}\leq$ $\sum_{l\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}(u_{2})_{l}\omega_{k-l}\|_{L^{2}L^{2}}$ (5.3)
+$\leq$ $\sum_{l\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}(u_{2})_{l}\|_{L^{2}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{\infty}L^{2}}$
+$\leq$ $\sum_{l\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\partial_{y}(u_{2})_{l}\|_{L^{2}L^{2}}^{\frac{1}{2}}\mathcal{E}_{k-l}$
+$\leq$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{E}_{l})^{\frac{1}{2}}(|l|\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{2}L^{2}})^{\frac{1}{2}}\mathcal{E}_{k-l}$
 $\leq$ $\sum_{l\in\mathbb{Z}}(|l|^{-\frac{1}{2}}\mathcal{E}_{l})^{\frac{1}{2}}(|l|^{-\frac{1}{2}}|l|\mathcal{E}_{l})^{\frac{1}{2}}\mathcal{E}_{k-l}\lesssim\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}$
 and
-$\|(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.9)
-$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}.$
+$\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}\leq$ $\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$ (5.4)
+$+\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}.$
 It remains to estimate the terms on the right-hand side of (5.4).
 For $l\not=0,k$, it holds that $|k|\lesssim|l||k-l|$, then
-$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|\omega_{k-l}\|_{L^{2}L^{2}}$
-$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}|l|^{-\frac{1}{2}}\mathcal{E}_{l}\nu^{-\frac{1}{6}}|k-l|^{-\frac{1}{3}}\mathcal{E}_{k-l}$
-$\lesssim$ $\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\mathcal{E}_{l}\mathcal{E}_{k-l}$
+$\sum_{l\in\mathbb{Z}\setminus\{0,k\}}$ $\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{\infty}L^{\infty}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}$
+$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\partial_{y}(u_{1})_{l}\|_{L^{\infty}L^{2}}^{\frac{1}{2}}$
+×$\times\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}^{\frac{5}{6}}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k-l}\|_{L^{2}L^{2}}^{\frac{1}{6}}$
+$\lesssim$ $\sum_{l\in\mathbb{Z}\setminus\{0,k\}}|l|^{-\frac{1}{2}}\mathcal{E}_{l}(\nu|k-l|)^{-\frac{1}{4}\cdot\frac{5}{6}}(\nu^{\frac{1}{2}}|k-l|)^{-\frac{1}{6}}\mathcal{E}_{k-l}$
+$\lesssim$ $\nu^{-\frac{7}{24}}|k|^{-\frac{3}{8}}\sum_{l\in\mathbb{Z}\setminus\{0,k\}}|l|^{-\frac{1}{8}}\mathcal{E}_{l}\mathcal{E}_{k-l}$
 and
-$\|\overline{u}_{1}\|_{L^{\infty}L^{\infty}}$ $\|\omega_{k}\|_{L^{2}L^{2}}+\|(u_{1})_{k}\|_{L^{2}L^{\infty}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
-$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}\|\omega_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\mathcal{E}_{0}\mathcal{E}_{k}.$
+[…]
+$\lesssim$ $\|\overline{\omega}\|_{L^{\infty}L^{2}}(\nu|k|)^{-\frac{1}{4}\cdot\frac{5}{6}}(\nu^{\frac{1}{2}}|k|)^{-\frac{1}{6}}\mathcal{E}_{k}$
+$+\|e^{c\nu^{\frac{1}{2}}t}(u_{1})_{k}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(\omega)_{k}\|_{L^{2}L^{2}}^{\frac{1}{2}}\|\overline{\omega}\|_{L^{\infty}L^{2}}$
+$\lesssim$ $\nu^{-\frac{7}{24}}|k|^{-\frac{3}{8}}\mathcal{E}_{0}\mathcal{E}_{k}.$
 Thus, we have
-$\|(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{1}{6}}|k|^{-\frac{1}{3}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}$ (5.10)
+$\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}\lesssim\nu^{-\frac{7}{24}}|k|^{-\frac{3}{8}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}$ (5.5)
 Combining (5.2), (5.3) and (5.5), we have
-$\mathcal{E}_{k}\lesssim$ $\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{3}}|k|^{\frac{1}{3}}\|(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|(f_{2})_{k}\|_{L^{2}L^{2}}$
-$\lesssim$ $\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l},$
+$\mathcal{E}_{k}\lesssim$ $\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}}$
+$\lesssim$ $\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{2}{3}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l},$
 $\mathcal{E}_{0}\lesssim$ $\|\overline{\omega}_{0}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{l\in\mathbb{Z}\setminus\{0\}}\mathcal{E}_{l}\mathcal{E}_{-l},$
 which gives that
-$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\lesssim\sum_{k\in\mathbb{Z}}\|\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{1}{2}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}.$
-Due to $\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{1/2}$, then one has
-$\sum_{k\in\mathbb{Z}}\|\omega_{0,k}\|_{L^{2}}\leq\epsilon_{1}C\nu^{\frac{1}{2}}.$
+$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\lesssim\sum_{k\in\mathbb{Z}}\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{2}{3}}\sum_{k\in\mathbb{Z}}\sum_{l\in\mathbb{Z}}\mathcal{E}_{l}\mathcal{E}_{k-l}.$
+Due to $\|u_{0}\|_{H^{\frac{7}{2}+}}\leq\epsilon_{1}\nu^{2/3}$ with
+$\sum_{k\in\mathbb{Z}}\|\Delta_{k}\omega_{0,k}\|_{L^{2}}\leq\epsilon_{1}C\nu^{\frac{2}{3}}.$
 If $\epsilon_{1}$ is suitably small, one can deduce from continuous argument that
-$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq\epsilon_{1}C\nu^{\frac{1}{2}}.$
+$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq\epsilon_{1}C\nu^{\frac{2}{3}}.$
 The proof is completed. ∎
@@
 S.
-Ding is supported by the Key Project of National Natural Science Foundation of China under Grant 12131010, National Natural Science Foundation of China under Grant 12271032 and Natural Science Foundation of Guangdong Province under Grant 2021A1515010249 and 2021A1515010303.
+Ding is supported by the Key Project of National Natural Science Foundation of China under Grant 12131010, National Natural Science Foundation of China under Grant 12271032.
 Z.